### v1

# AnomalyMatch: Discovering Rare Objects of Interest with Semi-supervised and Active Learning

Abstract
Anomaly detection in large datasets is essential in fields such as astronomy and computer vision; however, supervised methods typically require extensive anomaly labelling, which is often impractical. We present AnomalyMatch, an anomaly detection framework combining the semi-supervised FixMatch algorithm using EfficientNet classifiers with active learning. By treating anomaly detection as a semi-supervised binary classification problem, we efficiently utilise limited labelled and abundant unlabelled images. We allow iterative model refinement in a user interface for expert verification of high-confidence anomalies and correction of false positives. Built for astronomical data, AnomalyMatch generalises readily to other domains facing similar data challenges. Evaluations on the GalaxyMNIST astronomical dataset and the miniImageNet natural-image benchmark under severe class imbalance (1% anomalies for miniImageNet) display strong performance: starting from five to ten labelled anomalies and after three active learning cycles, we achieve an average AUROC of $0.95$ (miniImageNet) and $0.86$ (GalaxyMNIST), with respective AUPRC of $0.77$ and $0.71$. After active learning cycles, anomalies are ranked with 71% (miniImageNet) to 93% precision in the 1% of the highest-ranked images. AnomalyMatch is tailored for large-scale applications, efficiently processing predictions for $\approx$ 100 million images within three days on a single GPU. Integrated into ESA’s Datalabs platform, AnomalyMatch facilitates targeted discovery of scientifically valuable anomalies in vast astronomical datasets. Our results underscore the exceptional utility and scalability of this approach for anomaly discovery, highlighting the value of specialised approaches for domains characterised by severe label scarcity.
Keywords Semi-supervised learning  $\cdot$ Active learning  $\cdot$ Anomaly detection  $\cdot$ High-performance computing
1 Introduction
Identifying anomalies – rare and unusual outliers – in large datasets is a central challenge across machine learning and applied sciences  [1, 2]. Anomalies often hold critical relevance, either as scientifically significant discoveries or as artefacts requiring removal.
There are various anomaly detection approaches, but they generally fall into three categories based on availability of labelled, i.e. known, anomalies: fully supervised, semi-supervised, and unsupervised methods. Fully supervised anomaly detection typically achieves the best performance but demands substantial labelled anomaly examples, which are often unavailable or prohibitively expensive to obtain  [3]. Conversely, purely unsupervised methods operate without labels but can struggle to discriminate scientifically meaningful anomalies from irrelevant outliers or noise and do not easily allow to make use of known information and anomalies  [4]. Semi-supervised learning (SSL) approaches, which use unlabelled data along with some limited labelled instances, offer a practical balance, leveraging minimal supervision to improve performance significantly over fully unsupervised approaches [3, 5].
Real-world anomaly detection workflows often benefit from directly incorporating human expertise via active learning and user-in-the-loop approaches [6]. These frameworks allow iterative model refinement through user interaction, crucially exploiting expert domain knowledge to identify the most valuable anomalies from a scientific perspective. The Astronomaly tool [7] is a prominent example in astronomy, effectively combining active learning with anomaly detection to iteratively refine its predictions based on user feedback. Additionally, citizen science projects such as Galaxy Zoo [8] have demonstrated the potential for non-expert volunteers to accelerate and scale the labelling process. Our work embraces this philosophy, emphasizing active learning to guide the anomaly detection pipeline toward scientifically interesting anomalies, rather than merely statistically unusual ones.
In this work, we introduce AnomalyMatch, a semi-supervised anomaly detection tool that integrates FixMatch [5] with an EfficientNet-based neural architecture  [9], active learning, and a custom graphical user interface (GUI). This setup is then optimised for anomaly detection as a heavily imbalanced binary classification. Our framework enables domain experts or citizen scientists to iteratively label selected anomalies identified by the model during training, significantly improving detection performance through minimal additional labelling. Originally designed with astronomical datasets in mind, AnomalyMatch is highly generalizable to other domains with similar data characteristics. Moreover, our implementation prioritizes scalability: AnomalyMatch can efficiently process and analyse hundreds of millions of images on just one graphics card as demonstrated in the accompanying article  [10], making it particularly suitable for upcoming large-scale astronomical surveys and data repositories.
We validate AnomalyMatch using two datasets: miniImageNet (a natural-image benchmark) and GalaxyMNIST (an astronomy-specific galaxy morphology dataset). Our experiments quantify anomaly detection performance via area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC) metrics under severe class imbalance as well as a anomaly detection efficiency metric. These benchmark results clearly demonstrate the efficacy of our semi-supervised and active-learning-based approach. Furthermore, AnomalyMatch complements existing methods effectively; while frameworks like PyOD [11] or Astronomaly [7] primarily rely on classical methods (e.g., Isolation Forest), AnomalyMatch combines semi-supervised and active learning with training a neural network . Thus, our method excels in targeted anomaly discovery, empowering users to define and detect anomalies of specific scientific interest rather than general statistical outliers.
Our primary contributions are:
  • A novel semi-supervised anomaly detection pipeline (AnomalyMatch) combining FixMatch, EfficientNet, and active learning in a heavily optimised framework, effectively leveraging unlabelled data and greatly reducing labelling requirements.
  • A GUI-based active learning loop enabling human experts (or citizen scientists) to iteratively guide anomaly selection and improve model performance during training.
  • Comprehensive benchmarking of AnomalyMatch on both natural and astronomical image datasets, demonstrating excellent performance in terms of AUROC/AUPRC and anomaly detection efficiency with extremely limited labelled samples (five to ten labelled anomalies).
  • A highly efficient and scalable implementation capable of analysing 100 million images within days  [10], seamlessly integrated into ESA’s Datalabs platform.
  • A detailed discussion on implications for future large-scale anomaly detection applications, including potential expansions to new data modalities, advanced architectures (e.g., vision transformers), and explainable AI.
2 Related Work
2.1 Anomaly Detection Methods
Methods for anomaly detection, also known as outlier detection, are typically classified based on their dependence on labelled data into supervised, semi-supervised, and unsupervised learning paradigms [1].
Unsupervised methods identify anomalies based solely on deviations from the normal data distribution without labelled examples. Classical approaches include One-Class SVM [12], Isolation Forest [4], and Local Outlier Factor (LOF) [13], extensively available in libraries such as PyOD [11]. While effective on lower-dimensional or structured datasets, these methods often struggle with complex high-dimensional data like images, where feature extraction and similarity metrics are more challenging [14, 15].
Supervised methods, which require extensive labelled datasets containing anomalies, generally yield higher accuracy. However, fully labelled datasets are often impractical due to the difficulty and cost associated with obtaining labelled anomalies [3].
Semi-supervised methods, using a limited number of labelled examples and abundant unlabelled data, balance performance and practicality [16]. Deep learning-based semi-supervised anomaly detection approaches, such as Deep SAD [3], incorporate few labelled anomalies to guide learned representations. Similarly, Astronomaly [7] integrates active learning and semi-supervised learning principles into an astronomy-focused anomaly detection framework.
While classical approaches remain effective in simpler domains, neural networks have proven particularly promising for anomaly detection in complex, high-dimensional image data due to their inherent ability to learn rich, hierarchical representations from raw input [17].
2.2 Advances in Semi-Supervised Learning
General-purpose semi-supervised learning methods have rapidly advanced in recent years. Early SSL algorithms, including $\Pi$-Model, Pseudo-Label, VAT, Mean Teacher, and UDA, effectively leveraged unlabelled data through consistency regularization and pseudo-labelling [18, 19, 20]. Among these, FixMatch [5] has notably simplified the approach by combining consistency regularization and confidence-based pseudo-labelling. Specifically, FixMatch generates pseudo-labels from weakly augmented unlabelled data and strongly augmented versions and then trains the model to predict consistently on both, thus effectively making use of unlabelled data. This technique achieves state-of-the-art results on common benchmarks (e.g., CIFAR-10, SVHN) with minimal labelled examples.
Variants of FixMatch have been successfully applied to specific domains, such as remote sensing imagery [16], semantic segmentation [21, 22], and graph classification [23]. These adaptations highlight FixMatch’s flexibility in handling various data modalities and structured prediction tasks. The success of FixMatch and related approaches across diverse settings motivates their adaptation to anomaly detection tasks characterized by severe label scarcity and class imbalance. In this work, we leverage FixMatch by formulating anomaly detection as a two-class classification problem—“anomaly vs. normal”—thus fully exploiting the potential of SSL techniques in anomaly discovery.
Overall, our approach synthesizes state-of-the-art semi-supervised learning with domain-tailored active learning and scalable software implementation, bridging a methodological gap in anomaly detection efficacy, especially relevant to data-intensive scientific applications in domains such as astronomy.
3 Methods
Our anomaly detection system, AnomalyMatch, consists of a semi-supervised binary classifier (nominal vs. anomaly), using the FixMatch algorithm, combined with an active-learning loop and an intuitive guided user interface (GUI). We detail each component below, highlighting key design decisions.
Figure 1: Overview of the AnomalyMatch workflow. A FixMatch-based semi-supervised learning loop trains an EfficientNet backbone using weak and strong augmentations. An active learning interface supports user verification and labelling of additional samples. The final model can be applied in batch mode to large datasets, with detected anomalies exported for further analysis.
3.1 FixMatch for Anomaly Detection
We base our training methodology on FixMatch [5], initially proposed for multi-class semi-supervised learning. We adapt it specifically to anomaly detection, treating the problem as a binary classification: “nominal" (regular data) vs. “anomaly" (rare, interesting data) with heavy class imbalance. We use a small labelled set $\mathcal{D}_{L}=\{(x_{i},y_{i})\}$ and a large unlabelled set $\mathcal{D}_{U}=\{u_{j}\}$, typically with $|\mathcal{D}_{U}|\gg|\mathcal{D}_{L}|$ and severe imbalance in favour of nominal examples. A practical difference to applying FixMatch in our scenario is that additional labels for nominal data are easier to obtain, whereas anomaly labels are often extremely scarce or difficult to expand due to their rarity. To handle this severe imbalance, we include oversampling using a weighted random sampler with occurrence-based sampling probabilities of anomalies / nominal samples during training.
Let $f_{\theta}(x)$ denote the neural network model with parameters $\theta$, outputting the probability $p(y=\text{anomaly}|x)$ for input $x$. In each training step, unlabelled images from $\mathcal{D}_{U}$ undergo weak augmentation (simple flips, crops) to generate pseudo-labels if the model’s predicted probabilities exceed a high-confidence threshold $\tau$ (e.g., 0.95). A corresponding strongly augmented image (applying substantial random transformations) is then fed into the network, enforcing consistency by training the model to match its predictions to the pseudo-labels.
The overall loss comprises supervised binary cross-entropy loss (on labelled data) and unsupervised consistency loss (on confident pseudo-labelled data):
$L=L_{\text{sup}}+\lambda\cdot L_{\text{unsup}}$
We set use $\lambda=1$ to emphasize unlabelled data contributions. Importantly, our experiments showed that training for fewer iterations (on the order of hundreds of iterations rather than tens of epochs until full convergence) is beneficial in the human-in-the-loop setup, as prolonged training sessions risk overfitting to the initial labels. This is explored in more detail in the Results Section.
3.2 EfficientNet Backbone
For computational efficiency and high accuracy, we chose EfficientNetLite0 [9], which achieves efficiency by systematically balancing depth, width, and resolution of the network through compound scaling, as our backbone convolutional neural network using input images resized to $224\times 224$. Smaller input sizes, such as $150\times 150$, were also successfully employed in our companion paper searching Hubble Space Telescope images  [10]. . The network is initialized using ImageNet pre-trained weights, fine-tuning all layers during FixMatch semi-supervised training. Weight decay ($\ell_{2}$ regularization) is included to prevent overfitting.
3.3 Active Learning Loop and User Interface
Integrating human expertise is central to our pipeline. After initial training, the active learning loop identifies unlabelled images with the highest predicted anomaly scores, presenting them to the user through a dedicated interface (Figure 2). Users verify and label anomalies, correcting false positives and confirming interesting candidates, thus efficiently refining the model.
Figure 2: AnomalyMatch active learning interface built with ipywidgets. Users can inspect and label model-flagged candidates, adjust visual settings (e.g., brightness, contrast, RGB channels), and monitor performance via receiver operating characteristic and precision-recall curves. Model and label states can be saved and reloaded across sessions.
The GUI provides various practical features:
  • Image Processing Tools: Adjustments for brightness, contrast, unsharp masking, and channel-specific filters (RGB) to support precise labelling.
  • Performance Feedback: Immediate feedback through AUROC and AUPRC metrics if a held-out test set was defined initially.
  • Interactive Data Management: Users can load batches of unlabelled data, save session labels and models, and iteratively continue cycles of labelling and retraining.
This user-friendly interface leverages ipywidgets, facilitating interactive exploration within Jupyter notebooks. A prebuilt setup integrated within the European Space Agency’s (ESA) Datalabs platform further provides seamless access to large-scale astronomical datasets, such as the recent Euclid Q1 data release [24], enabling immediate practical use in scientific analyses.
3.4 Data Augmentations
Robust data augmentation is essential to effective semi-supervised learning. We started from augmentation strategies validated by MSMatch [16] and DistMSMatch [25] for the two-tiered augmentation approach:
  • Weak Augmentations: Minor transformations (horizontal flips, slight random crops) that preserve image content integrity, essential for accurate pseudo-label generation.
  • Strong Augmentations: Applying RandAugment [26], we perform aggressive transformations including rotations, colour distortions (brightness, contrast, solarization), sharpness variations, shearing, translations, and posterization. These diverse distortions encourage model robustness and invariance to image variability.
3.5 Implementation and Scalability
Our pipeline is built to ensure scalability, robustness, and reproducibility. Initially derived and extensively refactored from an existing PyTorch implementation of FixMatch [^1] https://github.com/LeeDoYup/FixMatch-pytorch — PyTorch code for FixMatch (Accessed 2025-03-28) , our architecture has been optimized for large-scale datasets, supporting efficient handling of datasets on terabyte scale (e.g. 100 million Hubble images) [10]. We handle data in HDF5 format to optimise performance during training and inference [27].
Practically important hyperparameters include the number of unlabelled images per training iteration (limited by GPU memory) and image input sizes, which directly influence training efficiency. Supported input formats include single-channel and RGB imagery (PNG, TIFF, JPEG).
Continuous integration and continuous deployment workflows incorporating over 70 unit tests ensure code robustness, reliability, and maintainability. This is essential given the complex interaction of semi-supervised and active learning components. Pending ESA licensing processes, the complete source code will be publicly available on GitHub  [^2] https://github.com/esa/AnomalyMatch, encouraging community collaboration and ongoing development.
4 Results
4.1 Datasets and Experimental Setup
We evaluate AnomalyMatch on two diverse datasets, one related to our primary application area in astronomy (GalaxyMNIST), and one standard benchmark commonly used in machine learning research (miniImageNet). MiniImageNet was specifically chosen as it is computationally more tractable than the full ImageNet dataset, making it well-suited for our experimentation.
GalaxyMNIST
[28] comprises annotated greyscale galaxy images derived from the Galaxy Zoo citizen science project [8]. Images originally have a resolution of $64\times 64$ pixels, which we upsampled to $224\times 224$ pixels for consistency across experiments. The dataset contains four morphological classes annotated by volunteers: Smooth Round, Smooth Cigar-shaped, Edge-on Disk, and Unbarred Spiral. For evaluation, we treated each morphological type sequentially as anomalies, using the other three classes as nominal samples.
miniImageNet
[29] is a widely used benchmark in few-shot learning. It consists of colour images grouped into 100 classes and is taken from the larger ImageNet dataset. Each image was resized from its original resolution to $224\times 224$ pixels. We deliberately selected five distinct object classes as anomalies to evaluate our model’s performance across diverse visual features: piano, hourglass, guitar, printer, and orange.
Figure 3 shows representative examples from both datasets.
Figure 3: Sample images from GalaxyMNIST (top) and MiniImageNet (bottom). Red borders indicate anomaly classes used during evaluation, while nominal samples are shown with white borders. For GalaxyMNIST all classes were tested as anomaly class in a dedicated run.
For GalaxyMNIST, we started experiments with a small labelled set of 40 images (10 anomalies, 30 nominal). For miniImageNet, initial labelled data included 500 images (five anomalies, 495 nominal). In both cases, each training step utilised an additional randomly sampled pool of 10,000 unlabelled images.
To simulate active learning, we performed three consecutive training cycles of 100 iterations each. After each cycle, the 10 highest-scoring anomalies and 10 highest-scoring false positives (nominal images with high anomaly classification scores by the model) were labelled and incorporated into the training set for the next cycle. The models are finally evaluated on all unlabelled images.
The FixMatch algorithm [5] formed the basis of our semi-supervised training procedure. Key hyperparameters included an exponential moving average (EMA) momentum of 0.99, confidence threshold ($\tau$) of 0.95, temperature of 0.5, unsupervised-to-supervised loss ratio of 1.0, a batch size of 16, learning rate of 0.0075 (with SGD optimizer, momentum 0.9), and weight decay of $7.5\times 10^{-4}$. We followed empirical results from [16, 25] in the selection of parameters.
Throughout the experiments, we employed three key metrics: Area Under the Receiver Operating Characteristic (AUROC), Area Under the Precision-Recall Curve (AUPRC), and Anomaly Detection Efficiency. AUROC and AUPRC provide complementary views of classification performance, especially important under severe class imbalance. Anomaly Detection Efficiency quantifies the fraction of anomalies correctly identified within a specified percentage of top-scoring samples, directly reflecting practical usefulness in scenarios involving human-in-the-loop validation and limited labelling resources.
[…]
Efficiency at $\text{Efficiency at }p\%=\frac{\text{Number of anomalies in top }p\%\text{ of scores}}{\text{Total number of anomalies}}\times 100\%$
where:
  • $\text{TPR}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FN}(t)}$ (true positive rate, also called recall)
  • $\text{FPR}(t)=\frac{\text{FP}(t)}{\text{FP}(t)+\text{TN}(t)}$ (false positive rate)
  • $\text{Precision}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FP}(t)}$
  • $\text{Recall}(t)=\text{TPR}(t)$
  • $\text{FPR}^{-1}(x)$ is the threshold value that gives an FPR of $x$
  • $\text{Recall}^{-1}(r)$ is the threshold value that gives a Recall of $r$
and $t\in[0,1]$ represents the classification threshold, and TP, FP, TN, FN denote counts of true positives, false positives, true negatives, and false negatives at threshold $t$.
4.2 miniImageNet Results
We summarize our experimental results on the miniImageNet dataset, focusing initially on the class hourglass as a representative example. Results for the other evaluated anomaly classes (printer, piano, guitar, and orange) are presented in Table 4.2.
Performance Metrics and Efficiency
We evaluate anomaly detection performance using the AUROC, AUPRC, and the anomaly detection efficiency metrics. Anomaly detection efficiency, a practical measure for real-world applications, quantifies the fraction of anomalies identified within the highest-scoring fraction of the dataset. Given the severe class imbalance (approximately 1% anomalies), high performance under this metric is critical to minimise manual inspection by domain experts.
Figure 4 shows representative receiver operation characteristic and Precision-Recall curves for the hourglass anomaly class after the final active learning cycle. Our approach achieved an AUROC of $0.96$ and AUPRC of $0.80$, indicating strong separation between nominal and anomalous samples.
Figure 4: Receiver operation characteristic (left) and Precision-Recall (right) curves after three active learning cycles for the miniImageNet ‘hourglass’ anomaly class starting from five labelled anomalies, adding ten after each iteration. The model achieves an AUROC of 0.96 and an AUPRC of 0.80, highlighting robust anomaly detection capability under severe class imbalance.
Figure 5 shows the anomaly detection efficiency curve after three iterations for the hourglass anomaly class. Remarkably, the method identifies 74.7% of anomalies within the top 1% highest-scoring images, demonstrating very high capability to retrieve relevant anomalies for the user. The top-scoring 0.1% data samples are all correctly identified anomalies. Successive active learning only marginally increase performance. A likely reason seems that learning gains from the initial sampling data, which remain the majority of labelled data throughout, can already be achieved within the first iteration.
(a) Efficiency progression across iterations: efficiency improves quickly, with strong performance already after the first iteration. The red dashed line represents perfect detection performance.
(b) Final performance after three iterations: 74.7% of all anomalies are found within the top 1% of the ranked predictions, with 10.3% already recovered in the top 0.1%.
Figure 5: Anomaly detection efficiency curves for the Hourglass class in miniImageNet. Each point on the x-axis represents a percentage of the ranked predictions inspected, sorted by descending anomaly score (i.e., the most anomalous images are checked first). The y-axis shows the percentage of true anomalies recovered within the inspected data subset. Active learning started with five labelled anomalies and adds ten samples after each iteration.
Table 4.2 summarizes performance metrics across all tested anomaly classes. The model consistently achieves high AUROC and AUPRC values, identifying approximately three-quarters of anomalies in the top 1% scoring images, and attaining precision between 65% and 81%. Average precisions for the top-scoring 0.1% of the data were 99.7%. This indicates strong practical utility for human reviewers who would want to quickly identify additional anomalies.
Table 1: Evaluation on unlabelled data for miniImageNet with different classes used as anomaly after three iterations with active learning. At 100% precision no more anomalies could have been found in the top-scoring 0.1% or 1% of the dataset, respectively. Training started from five labelled anomalies, adding ten after each iteration.
[…]
=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Anomaly Class AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
Guitar 0.95 0.76 10.24 100.00 72.32 70.19
Hourglass 0.96 0.80 10.24 100.00 74.72 72.52
Printer 0.98 0.86 10.24 100.00 81.44 79.04
Piano 0.95 0.75 10.24 100.00 71.36 69.25
Orange 0.92 0.65 10.24 100.00 64.16 62.27
Mean 0.95 $\pm$ 0.02 0.76 $\pm$ 0.08 10.24 $\pm$ 0.00 100.00 $\pm$ 0.00 72.80 $\pm$ 6.23 70.65 $\pm$ 6.05
In summary, these results demonstrate the effectiveness of our approach, highlighting its practical value by consistently ranking most anomalies highly and significantly reducing the labelling workload required from human experts for enabling such a result by starting from only five labelled samples.
4.3 GalaxyMNIST Results
For GalaxyMNIST, we illustrate our results using the Unbarred Spiral class as representative anomalies from the dataset. Notably, in this case the fraction of anomalies in the dataset (25%) is much higher compared to miniImageNet (1%).
Performance Metrics and Efficiency
Receiver operation characteristic and Precision-Recall curves (Figure 6) confirm the robustness of the model. Despite GalaxyMNIST’s classes being visually similar and challenging to distinguish even for human observers, the model achieves strong discrimination performance (AUROC: 0.889, AUPRC: 0.774), significantly above random chance.
Figure 6: Receiver operation characteristic (left) and Precision-Recall (right) curves for the Unbarred Spiral anomaly class from GalaxyMNIST. High AUROC and AUPRC indicate effective anomaly identification despite the inherent morphological similarities across galaxy classes. Training started from ten labelled anomalies, adding ten after each iteration.
Figure 7 shows how anomaly detection efficiency evolves across iterations. A clear improvement of the model is visible after each active learning step, indicating significantly enhanced anomaly ranking capabilities.
(a) Active learning efficiency across iterations: progressive improvements are visible over three iterations, with more anomalies recovered early in the ranked predictions.
(b) Final performance after three iterations: within the top 1% of the ranked predictions only anomalies are found
Figure 7: Anomaly detection efficiency curves for the Unbarred Spiral class in GalaxyMNIST. The x-axis shows the percentage of the data inspected, sorted by descending predicted anomaly score, while the y-axis reflects the cumulative percentage of true anomalies recovered. The model is trained using ten labelled anomalies initially, with ten more added after each active learning cycle.
The final anomaly detection efficiency curve in Figure 7 demonstrates strong final performance, particularly at lower inspection percentages. Remarkably, within the top 0.1% highest-scoring samples, 100% precision was achieved, thus successfully identifying anomalies immediately within a very small subset of the data. Successive active learning continue to increase performance in contrast to the experiments on MiniImageNet. One potential reason may lie in the overall greater number of anomalous samples in the dataset, leading to a potentially higher heterogeneity of anomalies.
Table 4.3 summarises results across all anomaly classes of GalaxyMNIST.
Table 2: Evaluation on unlabelled data for GalaxyMNIST with different classes used as anomaly after three iterations with active learning. At 100% precision no more anomalies could have been found in the top-scoring 0.1% or 1% of the dataset, respectively. Training started from ten labelled anomalies, adding ten after each iteration.
[…]
=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Anomaly Class AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
Smooth Round 0.97 0.91 0.36 100.00 3.97 98.99
Smooth Cigar-shaped 0.75 0.51 0.28 77.78 3.36 83.84
Edge-on Disk 0.85 0.66 0.36 100.00 3.72 92.93
Unbarred Spiral 0.89 0.77 0.36 100.00 3.93 97.98
Mean 0.86 $\pm$ 0.09 0.71 $\pm$ 0.17 0.34 $\pm$ 0.04 94.44 $\pm$ 11.11 3.75 $\pm$ 0.28 93.44 $\pm$ 6.92
Overall, anomaly detection performance on GalaxyMNIST varied between classes, reflecting visual and morphological similarities inherent in the dataset. Notably, three out of four classes achieved perfect anomaly detection precision within the top 0.1% of highest-scoring samples, demonstrating excellent efficiency in scenarios where quick anomaly identification from large data pools is required. Conversely, classes like Smooth Cigar-shaped and Edge-on Disk led to lower AUROC and AUPRC values, highlighting the inherent difficulty and potential ambiguity in visually separating these morphologies.
4.4 Ablation Studies
To evaluate the robustness and practical sensitivity of our approach to key training parameters, we conducted targeted ablation studies on the miniImageNet dataset, specifically focusing on the class guitar as anomalies. This class was picked as we achieved median performance on it.
4.4.1 Number of Initial Labels
We investigated the impact of varying the number of initial labels while maintaining the proportion of anomalies within these labels. Three scenarios were tested: 100 initial labels (one anomaly, 99 nominal), 500 initial labels (five anomalies, 495 nominal), and 1000 initial labels (10 anomalies, 990 nominal). After each active learning cycle, we correspondingly increased the number of actively labelled samples per iteration proportionally to the initial sample size: 10 additional labels per iteration for the smallest set, 20 labels for the intermediate set, and 40 labels for the largest set.
Detailed results for these ablations are summarised in Table 4.4.1.
Table 3: Effect of size of initial labelled dataset on anomaly detection performance in MiniImageNet. Each configuration starts with a different number of labelled samples and adds different number of labels per iteration. Despite large differences in initial label counts, performance remains robust across metrics, with perfect precision at the top 0.1% most anomalous samples for all scenarios. Beyond 500 (+20) labels no performance gains are observed.
[…]
=1.2mm {tabu} | X[0.45cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Labels AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
100 (1 / 99) + 10 per iteration 0.93 0.65 10.02 100.00 61.66 60.80
500 (5 / 495) + 20 per iteration 0.96 0.79 10.24 100.00 75.20 72.98
1000 (10 / 990) + 40 per iteration 0.95 0.78 10.50 100.00 75.00 70.42
Mean 0.95 $\pm$ 0.02 0.74 $\pm$ 0.08 10.25 $\pm$ 0.24 100.00 $\pm$ 0.00 70.62 $\pm$ 7.76 68.07 $\pm$ 6.42
Interestingly, we observed limited differences in performance metrics across all tested initial labelling conditions. AUROC and AUPRC values varied by less than $0.03$, and anomaly detection efficiency remained consistently high. Notably, all experiments achieved perfect precision at the stringent threshold of inspecting just the top 0.1% scoring data points, further underscoring the robustness and consistency of our active learning pipeline. This indicates our approach effectively utilises available labels and does not significantly benefit from substantially larger initial label sets, highlighting its practical efficacy even with minimal labelled data.
4.4.2 Training Iterations
To assess how training duration per cycle affects anomaly detection, we varied the number of FixMatch training iterations per cycle across four settings: 50, 100, 250, and 500. Results are presented in Table 4.4.2.
Table 4: Impact of FixMatch training duration per active learning step. Performance peaks at shorter training cycles (50–100 iterations), with diminishing returns and mild overfitting at longer durations. This supports our design choice of brief, iterative updates with rapid model adaptation.
[…]
=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Iterations AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
50 0.97 0.77 10.24 100.00 71.52 69.41
100 0.97 0.79 10.24 100.00 74.72 72.52
250 0.95 0.78 10.24 100.00 74.08 71.89
500 0.94 0.77 10.24 100.00 73.44 71.27
Mean 0.96 $\pm$ 0.02 0.78 $\pm$ 0.01 10.24 $\pm$ 0.00 100.00 $\pm$ 0.00 73.44 $\pm$ 1.38 71.27 $\pm$ 1.34
We observe that performance peaks at around 100 training iterations per cycle. Using more iterations results in slightly worse AUROC and AUPRC, which we attribute to overfitting on the initially provided labelled data. As the active learning loop dynamically updates the training set, excessively long training phases may entrench early biases. This supports our design choice of short training cycles, encouraging the model to adapt quickly to newly labelled examples.
4.4.3 Impact of Active Learning
We specifically investigated the impact of incorporating active learning by comparing our standard iterative labelling approach against two pure semi-supervised scenarios without additional human-in-the-loop labels, one with the same amount of initial labels and, respectively, same final amount (i.e. total number including those added during active learning). Figure 8 illustrates anomaly detection efficiency for those scenarios.
(a) Without active learning (same starting samples)
(b) Without active learning (same total samples)
(c) With active learning
(d) Comparison of approaches
Figure 8: Anomaly detection efficiency comparison with and without active learning. The actively learned model achieves notably better efficiency within the crucial 0.1–1% data inspection range for the same starting labels (compare (a) and (c)), indicating a meaningful advantage in human-in-the-loop anomaly searches. However, training with the full set of labels (as would be identified during active learning) from scratch leads to the best performance (b), highlighting the potential of re-using found anomalies in new semi-supervised runs. The direct comparison in (d) further illustrates the relative performance across all approaches, demonstrating how active learning bridges the gap between limited-label and full-label scenarios.
Table 5: Comparison of anomaly detection performance with and without active learning. Active learning at same starting labels yields higher AUPRC and precision within the top 1% most anomalous predictions. Using all actively gathered labels for semi-supervised training yields the highest performance, suggesting benefits from iterative human-in-the-loop feedback and re-use of confirmed anomalies.
[…]
=1.2mm {tabu} | X[0.8cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Active Learning AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
Active Learning 0.96 0.79 10.24 100.00 74.56 72.36
No Active Learning (same starting samples) 0.96 0.70 9.77 98.44 65.89 65.89
No Active Learning (same total samples) 0.99 0.85 10.41 100.00 78.70 75.16
Mean 0.97 $\pm$ 0.02 0.78 $\pm$ 0.08 10.14 $\pm$ 0.33 99.48 $\pm$ 0.90 73.05 $\pm$ 6.54 71.14 $\pm$ 4.75
While passive training (no additional active learning) yielded similar AUROC and AUPRC metrics, active learning exhibited significantly improved practical efficiency in regard to anomaly detection efficiency. Crucially, in the most relevant operational range (inspecting the top 0.1%–1% of scored samples), active learning consistently ranked more anomalies higher for the same starting samples. For instance, active learning allowed users to detect 75% of anomalies within the top 1% of the ranked dataset, while the passive model lagged behind in top-scoring predictions’ precision (particularly visible in the 0.1% inspection range).
This highlights a critical trade-off: despite marginally lower overall metric scores, the actively updated model significantly reduces human inspection efforts by focusing the user’s attention effectively, thus strongly supporting the use of active learning in real-world applications. However, retraining with all found anomalies still yields the best performance, likely due to a more balanced representation of anomaly samples during training.
4.5 Discussion
Our results highlight the practical efficiency and robustness of the AnomalyMatch framework, demonstrating significant potential to substantially reduce the manual review burden for anomaly identification tasks. The method efficiently leverages both computational resources and human labelling effort, achieving strong performance even with very limited labelled data. Crucially, our experiments reveal diminishing returns from significantly increasing the number of initial labels. This effect is consistent with earlier findings in semi-supervised learning theory and empirical studies that suggest performance gains saturate quickly beyond a modest label count [30]. This outcome underscores the method’s effective utilisation of limited supervision, suggesting that label sparsity — a common real-world constraint — is not a major limitation in practice.
The integration of active learning notably enhances anomaly detection efficiency, particularly in the critical inspection range (top 0.1%–1% of ranked data). While active learning’s impact on global performance metrics such as AUROC and AUPRC might appear moderate, its practical value in real-world scenarios cannot be overstated, as domain experts typically inspect only a small fraction of data. Thus, actively guided iterative labelling can substantially increase the productivity of domain experts, aligning well with real-world needs.
A key finding of our study is the pronounced risk of overfitting when extending training beyond the optimal number of iterations per active learning cycle. Notably, our experiments suggest the ideal balance is around 100 iterations per active learning cycle; longer cycles appear to reduce overall anomaly detection efficiency, highlighting the need for careful tuning and early stopping strategies. Such overfitting risk in SSL has been previously observed, especially in low-label regimes without strong regularisation [31]. Given that original FixMatch and its derivatives such as MSMatch typically employ extensive training iterations, our adaptation of shorter, frequent retraining cycles emerges as a beneficial modification tailored to active learning scenarios with heavily imbalanced classes and a simplified binary classification task.
Additionally, while we have demonstrated effectiveness across datasets with drastically different class imbalance characteristics –— extremely skewed (miniImageNet) and moderately balanced (GalaxyMNIST) —– further research is necessary to investigate the lower bounds of label numbers needed for robust anomaly detection performance. Preliminary findings from an accompanying paper focused on the Hubble Space Telescope dataset [10] suggest that AnomalyMatch might remain effective with even fewer initial labels, further reducing labelling costs. There, we also utilised a hybrid approach: initially using active learning to identify a sufficient number of anomalies, then transitioning to a fully semi-supervised regime for optimal performance.
4.6 Limitations
Despite strong results, our study acknowledges several important limitations that warrant further exploration:
Idealised Active Learning Assumptions:
Our experiments assumed an idealised active learning environment where human annotations are consistently accurate. In practical deployments, labelling errors and inconsistencies are likely, particularly in ambiguous cases. Additionally, our experiments assumed balanced incremental labelling (equal numbers of anomalies and nominal cases per active learning cycle), which may differ significantly in real-world usage where nominal labels are generally easier and cheaper to obtain than anomalies.
Pretraining Bias:
Performance on the miniImageNet dataset likely benefits from the EfficientNet backbone’s ImageNet pretraining. Such pretrained representations inherently capture generic features beneficial for image classification, possibly inflating performance metrics. Future studies should assess performance sensitivity when employing randomly initialised neural networks or representations pretrained on domain-specific data.
Limited Exploration of Class Imbalance Scenarios:
While we investigated datasets at two extremes of class imbalance (1% vs. 25% anomalies), real-world anomaly scenarios may present even more extreme imbalance, significantly affecting performance. Additional research is needed to characterise model behaviour across a broader spectrum of imbalance ratios.
Anomaly Homogeneity Assumption:
We simplified the anomaly detection task by treating all anomalies as belonging to a single implicit category distinct from nominal instances. This assumption might limit applicability in scenarios with highly heterogeneous anomalies, each requiring distinct learned representations. Incorporating methods explicitly capable of handling diverse anomaly categories could enhance generalisation.
Impact of Labelling Errors:
Our study does not examine the robustness of AnomalyMatch to labelling inaccuracies, an important consideration given the potential for human annotation errors during active learning. Exploring training strategies that explicitly handle or mitigate labelling noise is necessary for real-world robustness.
Addressing these limitations in future work will further strengthen the practical applicability of our approach, enabling its robust deployment in diverse, realistic scientific and industrial contexts.
5 Conclusion and Outlook
In this study, we introduced AnomalyMatch, a robust semi-supervised active learning framework tailored for efficiently identifying anomalies in large-scale image datasets. Our pipeline, combining FixMatch-based semi-supervised learning, EfficientNet as a computational backbone, and an interactive graphical user interface, demonstrated excellent performance across diverse datasets (GalaxyMNIST and miniImageNet), highlighting its capability to effectively operate under severe class imbalance and label scarcity—conditions frequently encountered in astronomy. By significantly reducing the manual review burden, AnomalyMatch offers substantial potential to streamline anomaly discovery processes, effectively harnessing minimal human input and maximising the efficiency of scarce domain-expert resources.
Looking ahead, several promising opportunities exist to further refine and extend the capabilities of AnomalyMatch. Incorporating explainable AI methods, such as attention maps, SHAP [32] values, or feature attribution techniques, could provide crucial interpretability and insight into model decision-making, increasing trust and facilitating model validation by domain experts. Additionally, extending the model to handle multimodal data—such as integrating image information with complementary spectral or time-series data—would allow richer contextual analysis, potentially enhancing anomaly detection accuracy in complex scientific scenarios. Further improvements could come from developing adaptive training strategies that dynamically calibrate training steps and parameters in response to real-time performance metrics, thus mitigating overfitting risks identified in our current approach. Finally, establishing curated, dedicated benchmark datasets with expert-validated labels, particularly in astronomy, would significantly support methodological advancements, comparison, and standardisation within the anomaly detection research community. Pursuing these enhancements will solidify AnomalyMatch as a versatile and impactful framework, beneficial not only to astronomy but broadly to scientific domains confronted with similar data challenges.
Acknowledgments
This work made use of AI tools for code generation and writing support. The authors would like to acknowledge valuable feedback received from Pedro Mas Buitrago, Laslo Erik Ruhberg and Maria Teresa Nardone.

### v2

@@
 Abstract
-Anomaly detection in large datasets is essential in fields such as astronomy and computer vision; however, supervised methods typically require extensive anomaly labelling, which is often impractical.
+Anomaly detection in large datasets is essential in astronomy and computer vision.
+However, due to a scarcity of labelled data, it is often infeasible to apply supervised methods to anomaly detection.
 We present AnomalyMatch, an anomaly detection framework combining the semi-supervised FixMatch algorithm using EfficientNet classifiers with active learning.
-By treating anomaly detection as a semi-supervised binary classification problem, we efficiently utilise limited labelled and abundant unlabelled images.
-We allow iterative model refinement in a user interface for expert verification of high-confidence anomalies and correction of false positives.
-Built for astronomical data, AnomalyMatch generalises readily to other domains facing similar data challenges.
-Evaluations on the GalaxyMNIST astronomical dataset and the miniImageNet natural-image benchmark under severe class imbalance (1% anomalies for miniImageNet) display strong performance: starting from five to ten labelled anomalies and after three active learning cycles, we achieve an average AUROC of $0.95$ (miniImageNet) and $0.86$ (GalaxyMNIST), with respective AUPRC of $0.77$ and $0.71$.
-After active learning cycles, anomalies are ranked with 71% (miniImageNet) to 93% precision in the 1% of the highest-ranked images.
-AnomalyMatch is tailored for large-scale applications, efficiently processing predictions for $\approx$ 100 million images within three days on a single GPU.
-Integrated into ESA’s Datalabs platform, AnomalyMatch facilitates targeted discovery of scientifically valuable anomalies in vast astronomical datasets.
+AnomalyMatch is tailored for large-scale applications and integrated into the ESA Datalabs science platform.
+In this method, we treat anomaly detection as a binary classification problem and efficiently utilise limited labelled and abundant unlabelled images for training.
+We enable active learning via a user interface for verification of high-confidence anomalies and correction of false positives.
+Evaluations on the GalaxyMNIST astronomical dataset and the miniImageNet natural-image benchmark under severe class imbalance display strong performance.
+Starting from five to ten labelled anomalies, we achieve an average AUROC of $0.96$ (miniImageNet) and $0.89$ (GalaxyMNIST), with respective AUPRC of $0.82$ and $0.77$.
+After three active learning cycles, anomalies are ranked with 76% (miniImageNet) to 94% (GalaxyMNIST) precision in the top 1% of the highest-ranking images by score.
+We compare to the established Astronomaly software on selected ’odd’ galaxies from the ’Galaxy Zoo - The Galaxy Challenge’ dataset, achieving comparable performance with an average AUROC of $0.83$.
 Our results underscore the exceptional utility and scalability of this approach for anomaly discovery, highlighting the value of specialised approaches for domains characterised by severe label scarcity.
-Keywords Semi-supervised learning  $\cdot$ Active learning  $\cdot$ Anomaly detection  $\cdot$ High-performance computing
+keywords:
+Semi-supervised learning – Active learning – Anomaly detection – High-performance computing
 1 Introduction
-Identifying anomalies – rare and unusual outliers – in large datasets is a central challenge across machine learning and applied sciences  [1, 2].
-Anomalies often hold critical relevance, either as scientifically significant discoveries or as artefacts requiring removal.
-There are various anomaly detection approaches, but they generally fall into three categories based on availability of labelled, i.e.
-known, anomalies: fully supervised, semi-supervised, and unsupervised methods.
-Fully supervised anomaly detection typically achieves the best performance but demands substantial labelled anomaly examples, which are often unavailable or prohibitively expensive to obtain  [3].
-Conversely, purely unsupervised methods operate without labels but can struggle to discriminate scientifically meaningful anomalies from irrelevant outliers or noise and do not easily allow to make use of known information and anomalies  [4].
-Semi-supervised learning (SSL) approaches, which use unlabelled data along with some limited labelled instances, offer a practical balance, leveraging minimal supervision to improve performance significantly over fully unsupervised approaches [3, 5].
-Real-world anomaly detection workflows often benefit from directly incorporating human expertise via active learning and user-in-the-loop approaches [6].
-These frameworks allow iterative model refinement through user interaction, crucially exploiting expert domain knowledge to identify the most valuable anomalies from a scientific perspective.
-The Astronomaly tool [7] is a prominent example in astronomy, effectively combining active learning with anomaly detection to iteratively refine its predictions based on user feedback.
-Additionally, citizen science projects such as Galaxy Zoo [8] have demonstrated the potential for non-expert volunteers to accelerate and scale the labelling process.
-Our work embraces this philosophy, emphasizing active learning to guide the anomaly detection pipeline toward scientifically interesting anomalies, rather than merely statistically unusual ones.
-In this work, we introduce AnomalyMatch, a semi-supervised anomaly detection tool that integrates FixMatch [5] with an EfficientNet-based neural architecture  [9], active learning, and a custom graphical user interface (GUI).
-This setup is then optimised for anomaly detection as a heavily imbalanced binary classification.
+Identifying anomalies – rare and unusual outliers – in large datasets is a central challenge in machine learning and the applied sciences  (Chandola et al., 2009; Pimentel et al., 2014).
+The term ‘anomaly’ has various definitions dependent upon the field they are being studied in.
+They can be telemetry outliers, instrumentation effects, objects exhibiting odd morphologies in images or outliers in well-measured distributions.
+Here we focus on searching for anomalies in image data.
+Some examples of astrophysical anomalies are gravitational lenses (Garvin et al., 2022; Euclid Collaboration et al., 2025), galaxy mergers (Pearson et al., 2022; O’Ryan et al., 2023), edge-on proto-planetary disks (Berghea et al., 2024), or ring galaxies (Abraham et al., 2025).
+Each of these objects provide unique insights into different aspects of astrophysics.
+Anomalies can hold critical relevance as scientifically meaningful discoveries, whether aiding our understanding in galaxy evolution, cosmology or the evolution of objects within the Galaxy.
+By their nature, anomalies are rare and require large or deep surveys to be detected in large quantities.
+Next-generation observatories, including Euclid (Laureijs et al., 2011) and the Vera C.
+Rubin Observatory (Ivezić et al., 2019), are poised to release unprecedented volumes of astronomical data across an unmatched area of the sky.
+With such large data volumes on the horizon, we require efficient methodologies to search these colossal datasets for rare objects where labelled data are limited or, potentially, non-existent.
+A key challenge in searching anomalies is that, given their rarity, there is typically a lack of labelled examples.
+Existing catalogues may consist of the order of tens of objects, for instance edge on protoplanetary disks (ALMA Partnership et al., 2015) or quadruply lensed quasars (Stern et al., 2021).
+We must also employ efficient techniques to identify them in datasets with the scale likely to contain them (i.e., sky spanning surveys such as Euclid).
+Due to the need for efficient processing of such datasets, machine learning methods are often employed for anomaly detection tasks.
+Supervised machine learning methods are most commonly applied to classification tasks in astronomy.
+Many works have successfully used supervised machine learning to classify astrophysical sources based on their morphology, and expand labelled sets of objects (Walmsley et al., 2023).
+However, they are often ineffective without large labelled datasets to train them.
+A promising avenue has been to fine-tune pre-trained neural network models for anomaly detection tasks.
+Fine-tuning involves taking a model pre-trained on a large dataset and continuing to train it on a new task or different data - e.g.
+classifying a specific anomaly.
+This reduces the need for large labelled datasets and training from scratch, but a few hundred examples are often still needed - larger than available catalogues for many types of anomaly - and effectiveness may vary depending on transferability of features and concept drift.
+Many works attempt to build the labelled datasets themselves to train supervised methods.
+Such approaches include tasking citizen scientists (e.g.
+Willett et al., 2013; Mantha et al., 2024) or have experts visually inspect a dataset to identify candidate anomalies (e.g.
+Acevedo Barroso et al., 2025).
+However, it has been found that special care must be taken when instructing individuals to search for different anomalies, or odd objects, as their definition can vary between different citizen scientists and experts (Mantha et al., 2024).
+Another issue is the size of observational datasets to search, as the scale of upcoming surveys makes visual inspection of a large portion of the data impossible.
+An approach which can bypass the need for labelled data entirely is that of unsupervised machine learning methods.
+These methods do not require labelled data at all, and utilise the structure of the underlying data itself in order to identify different classes of objects (Schölkopf et al., 2001; Zhao et al., 2019).
+While effective on lower-dimensional or structured datasets, these methods often struggle with complex high-dimensional data like images, where feature extraction and similarity metrics are more challenging (Goldstein & Uchida, 2016; Zimek et al., 2012).
+However, it is desirable to utilise the known anomalies, even if few, or annotated non-anomalous data when training a model.
+In many non-astronomical contexts, this has been achieved through semi-supervised learning techniques such as MSMatch (Gómez & Meoni, 2021), derived from the well-used FixMatch method (Sohn et al., 2020).
+These frameworks combine both supervised and unsupervised methods to learn about the data.
+MSMatch was able to achieve accuracy competitive with supervised methods on scene classification of satellite imagery with as few as five labelled examples per class while utilising thousands of unlabelled images.
+Another approach, combining classical feature space exploration with an element of user-in-the-loop active learning, is Astronomaly (Lochner & Bassett, 2021).
+This algorithm successfully leverages the features of galaxies to break down datasets into different classes, which can then be directed by a user to identify objects of interest.
+Astronomaly was successfully benchmarked on the Galaxy Zoo: The Galaxy Challenge dataset, a soft label dataset derived from the Galaxy Zoo 2 (Willett et al., 2013) campaign.
+They found they were able to recover ‘Odd’ galaxies from a larger dataset with no pre-existing labelled data of the objects.
+This work showed that real-world anomaly detection workflows often benefit from directly incorporating human expertise via active learning and user-in-the-loop approaches (Settles, 2009).
+Such a framework allows iterative model refinement through user interaction, crucially exploiting expert domain knowledge to identify the most valuable anomalies from a scientific perspective.
+Thus, combining a semi-supervised method - leveraging labelled and unlabelled data - with a user-in-the-loop workflow appears to be key for furthering anomaly detection methods in astronomy.
+In this work, we introduce AnomalyMatch - a method that embraces this philosophy.
+AnomalyMatch is a semi-supervised anomaly detection tool that integrates FixMatch with an EfficientNet-based neural network architecture  (Tan & Le, 2019), active learning, and a custom graphical user interface (GUI).
+This setup is optimised for anomaly detection as a heavily imbalanced binary classification.
 Our framework enables domain experts or citizen scientists to iteratively label selected anomalies identified by the model during training, significantly improving detection performance through minimal additional labelling.
-Originally designed with astronomical datasets in mind, AnomalyMatch is highly generalizable to other domains with similar data characteristics.
-Moreover, our implementation prioritizes scalability: AnomalyMatch can efficiently process and analyse hundreds of millions of images on just one graphics card as demonstrated in the accompanying article  [10], making it particularly suitable for upcoming large-scale astronomical surveys and data repositories.
-We validate AnomalyMatch using two datasets: miniImageNet (a natural-image benchmark) and GalaxyMNIST (an astronomy-specific galaxy morphology dataset).
-Our experiments quantify anomaly detection performance via area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC) metrics under severe class imbalance as well as a anomaly detection efficiency metric.
-These benchmark results clearly demonstrate the efficacy of our semi-supervised and active-learning-based approach.
-Furthermore, AnomalyMatch complements existing methods effectively; while frameworks like PyOD [11] or Astronomaly [7] primarily rely on classical methods (e.g., Isolation Forest), AnomalyMatch combines semi-supervised and active learning with training a neural network .
-Thus, our method excels in targeted anomaly discovery, empowering users to define and detect anomalies of specific scientific interest rather than general statistical outliers.
+Originally designed with astronomical datasets in mind, AnomalyMatch is highly generalisable to other domains with similar data characteristics.
+Moreover, our implementation prioritises scalability: AnomalyMatch can efficiently process and analyse hundreds of millions of images on just one graphics card, as demonstrated in the accompanying article (O’Ryan & Gómez, 2025), which makes it particularly suitable for upcoming large-scale astronomical surveys and data repositories.
+Thus, our method excels in targeted anomaly discovery, empowering users to define and detect anomalies of scientific interest rather than general statistical outliers.
 Our primary contributions are:
-  • A novel semi-supervised anomaly detection pipeline (AnomalyMatch) combining FixMatch, EfficientNet, and active learning in a heavily optimised framework, effectively leveraging unlabelled data and greatly reducing labelling requirements.
-  • A GUI-based active learning loop enabling human experts (or citizen scientists) to iteratively guide anomaly selection and improve model performance during training.
-  • Comprehensive benchmarking of AnomalyMatch on both natural and astronomical image datasets, demonstrating excellent performance in terms of AUROC/AUPRC and anomaly detection efficiency with extremely limited labelled samples (five to ten labelled anomalies).
-  • A highly efficient and scalable implementation capable of analysing 100 million images within days  [10], seamlessly integrated into ESA’s Datalabs platform.
-  • A detailed discussion on implications for future large-scale anomaly detection applications, including potential expansions to new data modalities, advanced architectures (e.g., vision transformers), and explainable AI.
-2 Related Work
-2.1 Anomaly Detection Methods
-Methods for anomaly detection, also known as outlier detection, are typically classified based on their dependence on labelled data into supervised, semi-supervised, and unsupervised learning paradigms [1].
-Unsupervised methods identify anomalies based solely on deviations from the normal data distribution without labelled examples.
-Classical approaches include One-Class SVM [12], Isolation Forest [4], and Local Outlier Factor (LOF) [13], extensively available in libraries such as PyOD [11].
-While effective on lower-dimensional or structured datasets, these methods often struggle with complex high-dimensional data like images, where feature extraction and similarity metrics are more challenging [14, 15].
-Supervised methods, which require extensive labelled datasets containing anomalies, generally yield higher accuracy.
-However, fully labelled datasets are often impractical due to the difficulty and cost associated with obtaining labelled anomalies [3].
-Semi-supervised methods, using a limited number of labelled examples and abundant unlabelled data, balance performance and practicality [16].
-Deep learning-based semi-supervised anomaly detection approaches, such as Deep SAD [3], incorporate few labelled anomalies to guide learned representations.
-Similarly, Astronomaly [7] integrates active learning and semi-supervised learning principles into an astronomy-focused anomaly detection framework.
-While classical approaches remain effective in simpler domains, neural networks have proven particularly promising for anomaly detection in complex, high-dimensional image data due to their inherent ability to learn rich, hierarchical representations from raw input [17].
-2.2 Advances in Semi-Supervised Learning
-General-purpose semi-supervised learning methods have rapidly advanced in recent years.
-Early SSL algorithms, including $\Pi$-Model, Pseudo-Label, VAT, Mean Teacher, and UDA, effectively leveraged unlabelled data through consistency regularization and pseudo-labelling [18, 19, 20].
-Among these, FixMatch [5] has notably simplified the approach by combining consistency regularization and confidence-based pseudo-labelling.
-Specifically, FixMatch generates pseudo-labels from weakly augmented unlabelled data and strongly augmented versions and then trains the model to predict consistently on both, thus effectively making use of unlabelled data.
-This technique achieves state-of-the-art results on common benchmarks (e.g., CIFAR-10, SVHN) with minimal labelled examples.
-Variants of FixMatch have been successfully applied to specific domains, such as remote sensing imagery [16], semantic segmentation [21, 22], and graph classification [23].
-These adaptations highlight FixMatch’s flexibility in handling various data modalities and structured prediction tasks.
-The success of FixMatch and related approaches across diverse settings motivates their adaptation to anomaly detection tasks characterized by severe label scarcity and class imbalance.
-In this work, we leverage FixMatch by formulating anomaly detection as a two-class classification problem—“anomaly vs.
-normal”—thus fully exploiting the potential of SSL techniques in anomaly discovery.
-Overall, our approach synthesizes state-of-the-art semi-supervised learning with domain-tailored active learning and scalable software implementation, bridging a methodological gap in anomaly detection efficacy, especially relevant to data-intensive scientific applications in domains such as astronomy.
+  • A novel semi-supervised anomaly detection pipeline (AnomalyMatch) combining FixMatch streamlined for heavy class imbalance, EfficientNet neural networks, and active learning in a heavily optimised framework, effectively leveraging unlabelled data and greatly reducing labelling requirements;
+  • A GUI-based active learning loop enabling experts to iteratively guide anomaly selection and improve model performance during training;
+  • Seamless integration of established astronomy image processing tools, such as different normalisation methods from Astropy (Astropy Collaboration et al., 2013, 2018) and FITS file support.
+  • Comprehensive benchmarking of AnomalyMatch on both daily-life and astronomical image datasets, demonstrating excellent performance in terms of Area Under the Receiver Operating Characteristic (AUROC), Area Under the Precision-Recall Curve (AUPRC), and Anomaly Detection Efficiency with extremely limited labelled samples (five to ten labelled anomalies);
+  • A highly efficient and scalable implementation capable of analysing 100 million images within days  (O’Ryan & Gómez, 2025), seamlessly integrated into the European Space Agency’s (ESA) Datalabs platform [^1] https://datalabs.esa.int/ (Navarro et al., 2024);
+  • A detailed discussion on implications for future large-scale anomaly detection applications, including potential expansions to new data modalities, advanced architectures (e.g.
+vision transformers), and explainable AI.
+Section 2 presents the components of the AnomalyMatch pipeline and provides the reasoning behind its key design choices.
+Section 3 describes the extensive testing and benchmarking of AnomalyMatch on multiple datasets and compares directly to the well used anomaly detection software Astronomaly on a dataset of anomalies from Galaxy Zoo 2.
+Section 4 discusses the significance of our anomaly detection framework and outlines its assumptions and limitations.
+Section 5 includes concluding remarks and discusses potential avenues for future work.
 2 Methods
-Our anomaly detection system, AnomalyMatch, consists of a semi-supervised binary classifier (nominal vs.
-anomaly), using the FixMatch algorithm, combined with an active-learning loop and an intuitive guided user interface (GUI).
-We detail each component below, highlighting key design decisions.
+Our anomaly detection method, AnomalyMatch, consists of a semi-supervised binary classifier (normal vs.
+anomaly), using the FixMatch algorithm.
+We have combined this with an active-learning loop and an intuitive GUI.
+We detail each component here, highlighting key design decisions.
 Figure 1: Overview of the AnomalyMatch workflow.
@@
 2.1 FixMatch for Anomaly Detection
-We base our training methodology on FixMatch [5], initially proposed for multi-class semi-supervised learning.
-We adapt it specifically to anomaly detection, treating the problem as a binary classification: “nominal" (regular data) vs.
-“anomaly" (rare, interesting data) with heavy class imbalance.
-We use a small labelled set $\mathcal{D}_{L}=\{(x_{i},y_{i})\}$ and a large unlabelled set $\mathcal{D}_{U}=\{u_{j}\}$, typically with $|\mathcal{D}_{U}|\gg|\mathcal{D}_{L}|$ and severe imbalance in favour of nominal examples.
-A practical difference to applying FixMatch in our scenario is that additional labels for nominal data are easier to obtain, whereas anomaly labels are often extremely scarce or difficult to expand due to their rarity.
-To handle this severe imbalance, we include oversampling using a weighted random sampler with occurrence-based sampling probabilities of anomalies / nominal samples during training.
-Let $f_{\theta}(x)$ denote the neural network model with parameters $\theta$, outputting the probability $p(y=\text{anomaly}|x)$ for input $x$.
+We base our training methodology on FixMatch (Sohn et al., 2020), initially proposed for multi-class semi-supervised learning.
+FixMatch generates pseudo-labels from weakly and strongly augmented unlabelled data and then trains the model to predict consistently on both.
+Variants of FixMatch have been successfully applied to specific domains, such as remote sensing imagery (Gómez & Meoni, 2021), semantic segmentation (Martí i Rabadán et al., 2022; Yang et al., 2022), and graph classification (Jalali et al., 2024).
+These adaptations highlight FixMatch’s flexibility in handling various data types and prediction tasks.
+The success of FixMatch and related approaches across diverse settings motivates their adaptation to anomaly detection tasks characterised by severe label scarcity and class imbalance.
+We adapt FixMatch to treat the problem of anomaly detection as a binary classification: "normal" (regular data) vs.
+"anomaly" (rare, interesting data) with heavy class imbalance.
+We use a labelled set $\mathcal{D}_{L}=\{(x_{i},y_{i})\}$ and an unlabelled set $\mathcal{D}_{U}=\{u_{j}\}$, typically with $|\mathcal{D}_{U}|\gg|\mathcal{D}_{L}|$ and severe imbalance in favour of normal examples.
+A practical difference to applying FixMatch in our scenario is that additional labels for normal data are easier to obtain, whereas the anomaly labelled dataset is small and difficult to expand due to their rarity.
+To handle this severe imbalance, we include oversampling using a weighted random sampler with occurrence-based sampling probabilities of anomalies / normal samples during training.
+To train in an unsupervised manner, consistency regularisation is used.
+Let $f_{\theta}(x)$ denote the neural network model with parameters $\theta$, outputting the score $p(y=\text{anomaly}|x)$ for input $x$.
 In each training step, unlabelled images from $\mathcal{D}_{U}$ undergo weak augmentation (simple flips, crops) to generate pseudo-labels if the model’s predicted probabilities exceed a high-confidence threshold $\tau$ (e.g., 0.95).
-A corresponding strongly augmented image (applying substantial random transformations) is then fed into the network, enforcing consistency by training the model to match its predictions to the pseudo-labels.
-The overall loss comprises supervised binary cross-entropy loss (on labelled data) and unsupervised consistency loss (on confident pseudo-labelled data):
-$L=L_{\text{sup}}+\lambda\cdot L_{\text{unsup}}$
-We set use $\lambda=1$ to emphasize unlabelled data contributions.
-Importantly, our experiments showed that training for fewer iterations (on the order of hundreds of iterations rather than tens of epochs until full convergence) is beneficial in the human-in-the-loop setup, as prolonged training sessions risk overfitting to the initial labels.
-This is explored in more detail in the Results Section.
+The same unlabelled images are strongly augmented (applying substantial random transformations) and then fed into the network, enforcing consistency by training the model to match its predictions to the pseudo-labels.
+This only comprises half the loss function when training the model.
+The overall loss combines the unsupervised consistency regularisation with a supervised binary cross-entropy loss calculated using labelled data as
+$L=L_{\text{sup}}+\lambda\cdot L_{\text{unsup}}.$ (1)
+We set $\lambda=1$ in Equation 1 to emphasize unlabelled data contributions.
+Importantly, we find that training for fewer iterations (on the order of hundreds of iterations rather than tens of epochs) is beneficial in the human-in-the-loop setup, as prolonged training sessions risk overfitting to the initial labels.
+This is explored in more detail in the Section 3.
 2.2 EfficientNet Backbone
-For computational efficiency and high accuracy, we chose EfficientNetLite0 [9], which achieves efficiency by systematically balancing depth, width, and resolution of the network through compound scaling, as our backbone convolutional neural network using input images resized to $224\times 224$.
-Smaller input sizes, such as $150\times 150$, were also successfully employed in our companion paper searching Hubble Space Telescope images  [10].
-.
-The network is initialized using ImageNet pre-trained weights, fine-tuning all layers during FixMatch semi-supervised training.
-Weight decay ($\ell_{2}$ regularization) is included to prevent overfitting.
+For computational efficiency and high accuracy, we chose EfficientNetLite0 (Tan & Le, 2019).
+This architecture achieves improved efficiency by systematically balancing depth, width, and resolution of the network through compound scaling as our backbone convolutional neural network uses input images resized to $224\times 224$.
+Smaller input sizes, such as $150\times 150$, were also successfully employed in our companion paper searching Hubble Space Telescope images  (O’Ryan & Gómez, 2025).
+The network is initialised using ImageNet pre-trained weights, fine-tuning all layers during FixMatch semi-supervised training.
+Weight decay ($\ell_{2}$ regularisation) is included to prevent overfitting.
 2.3 Active Learning Loop and User Interface
-Integrating human expertise is central to our pipeline.
-After initial training, the active learning loop identifies unlabelled images with the highest predicted anomaly scores, presenting them to the user through a dedicated interface (Figure 2).
-Users verify and label anomalies, correcting false positives and confirming interesting candidates, thus efficiently refining the model.
+Leveraging human expertise is central to our pipeline.
+After initial training, the unlabelled data used is scored and ranked from most to least anomalous.
+A user is able to then confirm or correct the algorithm, both expanding the training set and refining the model.
+This also gives initial feedback to the user on the models performance.
+Figure 2 shows the dedicated interface through which the user conducts the active learning loop.
 Figure 2: AnomalyMatch active learning interface built with ipywidgets.
-Users can inspect and label model-flagged candidates, adjust visual settings (e.g., brightness, contrast, RGB channels), and monitor performance via receiver operating characteristic and precision-recall curves.
+Users can inspect and label model-flagged candidates, adjust visual settings (e.g., brightness, contrast, RGB channels), and monitor performance via AUROC and AUPRC curves.
 Model and label states can be saved and reloaded across sessions.
 The GUI provides various practical features:
-  • Image Processing Tools: Adjustments for brightness, contrast, unsharp masking, and channel-specific filters (RGB) to support precise labelling.
-  • Performance Feedback: Immediate feedback through AUROC and AUPRC metrics if a held-out test set was defined initially.
-  • Interactive Data Management: Users can load batches of unlabelled data, save session labels and models, and iteratively continue cycles of labelling and retraining.
+  • Image Display Tools. Adjustments for brightness, contrast, unsharp masking, and channel-specific filters (RGB) to support precise labelling;
+  • Image Processing Tools. Adjustments for the normalisation and stretch applied when loading, training and evaluating images;
+  • Performance Feedback. Immediate feedback through AUROC and AUPRC metrics if a test set was defined initially;
+  • Interactive Data Management. Users can load batches of unlabelled data, save session labels and models, and iteratively continue cycles of labelling and retraining.
 This user-friendly interface leverages ipywidgets, facilitating interactive exploration within Jupyter notebooks.
-A prebuilt setup integrated within the European Space Agency’s (ESA) Datalabs platform further provides seamless access to large-scale astronomical datasets, such as the recent Euclid Q1 data release [24], enabling immediate practical use in scientific analyses.
-3.4 Data Augmentations
+A prebuilt setup integrated within ESA Datalabs further provides seamless access to large-scale astronomical datasets, such as the Euclid Quick Data Release (Q1; Aussel et al., 2025).
+This enables immediate practical use in scientific analyses.
+2.4 Data Processing and Augmentations
 Robust data augmentation is essential to effective semi-supervised learning.
-We started from augmentation strategies validated by MSMatch [16] and DistMSMatch [25] for the two-tiered augmentation approach:
-  • Weak Augmentations: Minor transformations (horizontal flips, slight random crops) that preserve image content integrity, essential for accurate pseudo-label generation.
-  • Strong Augmentations: Applying RandAugment [26], we perform aggressive transformations including rotations, colour distortions (brightness, contrast, solarization), sharpness variations, shearing, translations, and posterization.
+We start from augmentation strategies validated by MSMatch (Gómez & Meoni, 2021) and DistMSMatch (Östman et al., 2023) for the two-tiered augmentation approach:
+  • Weak Augmentations. Minor transformations (horizontal flips, slight random crops) that preserve image content integrity, essential for accurate pseudo-label generation;
+  • Strong Augmentations.
+Applying RandAugment (Cubuk et al., 2020), we perform aggressive transformations including rotations, colour distortions (brightness, contrast, solarisation), sharpness variations, shearing, translations, and posterisation.
 These diverse distortions encourage model robustness and invariance to image variability.
@@
 Our pipeline is built to ensure scalability, robustness, and reproducibility.
-Initially derived and extensively refactored from an existing PyTorch implementation of FixMatch [^1] https://github.com/LeeDoYup/FixMatch-pytorch — PyTorch code for FixMatch (Accessed 2025-03-28) , our architecture has been optimized for large-scale datasets, supporting efficient handling of datasets on terabyte scale (e.g.
-100 million Hubble images) [10].
-We handle data in HDF5 format to optimise performance during training and inference [27].
-Practically important hyperparameters include the number of unlabelled images per training iteration (limited by GPU memory) and image input sizes, which directly influence training efficiency.
-Supported input formats include single-channel and RGB imagery (PNG, TIFF, JPEG).
-Continuous integration and continuous deployment workflows incorporating over 70 unit tests ensure code robustness, reliability, and maintainability.
+Initially derived and extensively refactored from an existing PyTorch implementation of FixMatch [^2] https://github.com/LeeDoYup/FixMatch-pytorch – PyTorch code for FixMatch (Accessed 2025-03-28), our architecture has been optimised for large-scale datasets, supporting efficient handling of datasets on terabyte scale (e.g.
+100 million Hubble images O’Ryan & Gómez, 2025).
+We handle data in HDF5 or Zarr format to optimise performance during training and inference (Folk et al., 2011).
+Important hyperparameters include the number of unlabelled images per training cycle (limited by GPU memory) and image input sizes, which directly influence training efficiency.
+Supported input formats include single-channel and RGB imagery (PNG, TIFF, JPEG) and FITS files.
+In the case of FITS files AnomalyMatch can specify one or 3 extensions to use.
+Furthermore, the user can select between different normalisation options such as the simple conversion, logarithmic or Asinh stretching or a Z-Scale based linear stretch.
+Continuous integration and deployment workflows incorporating over 70 unit tests ensure code robustness, reliability, and maintainability.
 This is essential given the complex interaction of semi-supervised and active learning components.
-Pending ESA licensing processes, the complete source code will be publicly available on GitHub  [^2] https://github.com/esa/AnomalyMatch, encouraging community collaboration and ongoing development.
+The complete source code of AnomalyMatch is publicly available on GitHub [^3] https://github.com/esa/AnomalyMatch, encouraging community collaboration and ongoing development.
 3 Results
-4.1 Datasets and Experimental Setup
-We evaluate AnomalyMatch on two diverse datasets, one related to our primary application area in astronomy (GalaxyMNIST), and one standard benchmark commonly used in machine learning research (miniImageNet).
-MiniImageNet was specifically chosen as it is computationally more tractable than the full ImageNet dataset, making it well-suited for our experimentation.
-GalaxyMNIST
-[28] comprises annotated greyscale galaxy images derived from the Galaxy Zoo citizen science project [8].
-Images originally have a resolution of $64\times 64$ pixels, which we upsampled to $224\times 224$ pixels for consistency across experiments.
-The dataset contains four morphological classes annotated by volunteers: Smooth Round, Smooth Cigar-shaped, Edge-on Disk, and Unbarred Spiral.
-For evaluation, we treated each morphological type sequentially as anomalies, using the other three classes as nominal samples.
-miniImageNet
-[29] is a widely used benchmark in few-shot learning.
+3.1 Experimental Setup
+We evaluate AnomalyMatch on three datasets: miniImageNet (Vinyals et al., 2016), a common benchmark used in machine learning research; GalaxyMNIST (Walmsley et al., 2022), a dataset pertinent to our primary application area, astronomy; and a subset of $\approx 60,000$ galaxy images extracted from Galaxy Zoo 2 (Willett et al., 2013) which were used for the ’Galaxy Challenge’ on Kaggle [^4] https://www.kaggle.com/c/galaxy-zoo-the-galaxy-challenge.
+The initial configurations (label counts and anomaly ratios) varied by dataset.
+Detailed descriptions of each dataset, along with their respective initial configurations, are provided in the following Sections.
+Figure 3 shows representative examples from both datasets.
+Figure 3: Sample images from GalaxyMNIST (top), Galaxy Zoo 2 (centre), and MiniImageNet (bottom).
+Red borders indicate anomaly classes used during evaluation, while normal samples are shown with white borders.
+For GalaxyMNIST all classes were tested as anomaly class in a dedicated run.
+The FixMatch algorithm (Sohn et al., 2020) formed the basis of our semi-supervised training procedure.
+Key hyperparameters included an exponential moving average (EMA) momentum of 0.99, confidence threshold ($\tau$) of 0.95, temperature of 0.5, unsupervised-to-supervised loss ratio ($\lambda$ in Equation 1) of 1.0, a batch size of 16, learning rate of 0.0075 (with a stochastic gradient descent (SGD) optimiser, momentum 0.9), and weight decay of $7.5\times 10^{-4}$.
+We followed empirical results from Gómez & Meoni (2021); Östman et al.
+(2023) in the selection of parameters.
+To simulate active learning for all three datasets we performed three consecutive training cycles of 100 iterations each.
+At the end of each cycle, the 10 highest-scoring anomalies and 10 highest-scoring false positives (normal images with high anomaly classification scores by the model) were labelled and incorporated into the training set for the next cycle.
+The models were finally evaluated on all unlabelled images.
+Throughout the experiments, we employed three key performance metrics: AUROC, AUPRC, and Anomaly Detection Efficiency.
+AUROC and AUPRC provide complementary views of classification performance, especially important under severe class imbalance.
+The Anomaly Detection Efficiency quantifies the fraction of anomalies correctly identified within a specified percentage of top-scoring samples, directly reflecting practical usefulness in scenarios involving human-in-the-loop validation and limited labelling resources.
+The mathematical definitions of all three metrics are given in Appendix A.
+3.2 miniImageNet
+miniImageNet (Vinyals et al., 2016) is a widely used benchmark in few-shot learning.
 It consists of colour images grouped into 100 classes and is taken from the larger ImageNet dataset.
+It was specifically chosen as it is computationally more efficient than the full ImageNet dataset, making it well-suited for our experimentation.
 Each image was resized from its original resolution to $224\times 224$ pixels.
-We deliberately selected five distinct object classes as anomalies to evaluate our model’s performance across diverse visual features: piano, hourglass, guitar, printer, and orange.
-Figure 3 shows representative examples from both datasets.
-Figure 3: Sample images from GalaxyMNIST (top) and MiniImageNet (bottom).
-Red borders indicate anomaly classes used during evaluation, while nominal samples are shown with white borders.
-For GalaxyMNIST all classes were tested as anomaly class in a dedicated run.
-For GalaxyMNIST, we started experiments with a small labelled set of 40 images (10 anomalies, 30 nominal).
-For miniImageNet, initial labelled data included 500 images (five anomalies, 495 nominal).
-In both cases, each training step utilised an additional randomly sampled pool of 10,000 unlabelled images.
-To simulate active learning, we performed three consecutive training cycles of 100 iterations each.
-After each cycle, the 10 highest-scoring anomalies and 10 highest-scoring false positives (nominal images with high anomaly classification scores by the model) were labelled and incorporated into the training set for the next cycle.
-The models are finally evaluated on all unlabelled images.
-The FixMatch algorithm [5] formed the basis of our semi-supervised training procedure.
-Key hyperparameters included an exponential moving average (EMA) momentum of 0.99, confidence threshold ($\tau$) of 0.95, temperature of 0.5, unsupervised-to-supervised loss ratio of 1.0, a batch size of 16, learning rate of 0.0075 (with SGD optimizer, momentum 0.9), and weight decay of $7.5\times 10^{-4}$.
-We followed empirical results from [16, 25] in the selection of parameters.
-Throughout the experiments, we employed three key metrics: Area Under the Receiver Operating Characteristic (AUROC), Area Under the Precision-Recall Curve (AUPRC), and Anomaly Detection Efficiency.
-AUROC and AUPRC provide complementary views of classification performance, especially important under severe class imbalance.
-Anomaly Detection Efficiency quantifies the fraction of anomalies correctly identified within a specified percentage of top-scoring samples, directly reflecting practical usefulness in scenarios involving human-in-the-loop validation and limited labelling resources.
-[…]
-[…]
-Efficiency at $\text{Efficiency at }p\%=\frac{\text{Number of anomalies in top }p\%\text{ of scores}}{\text{Total number of anomalies}}\times 100\%$
-where:
-  • $\text{TPR}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FN}(t)}$ (true positive rate, also called recall)
-  • $\text{FPR}(t)=\frac{\text{FP}(t)}{\text{FP}(t)+\text{TN}(t)}$ (false positive rate)
-  • $\text{Precision}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FP}(t)}$
-  • $\text{Recall}(t)=\text{TPR}(t)$
-  • $\text{FPR}^{-1}(x)$ is the threshold value that gives an FPR of $x$
-  • $\text{Recall}^{-1}(r)$ is the threshold value that gives a Recall of $r$
-and $t\in[0,1]$ represents the classification threshold, and TP, FP, TN, FN denote counts of true positives, false positives, true negatives, and false negatives at threshold $t$.
-4.2 miniImageNet Results
-We summarize our experimental results on the miniImageNet dataset, focusing initially on the class hourglass as a representative example.
-Results for the other evaluated anomaly classes (printer, piano, guitar, and orange) are presented in Table 4.2.
-Performance Metrics and Efficiency
-We evaluate anomaly detection performance using the AUROC, AUPRC, and the anomaly detection efficiency metrics.
-Anomaly detection efficiency, a practical measure for real-world applications, quantifies the fraction of anomalies identified within the highest-scoring fraction of the dataset.
-Given the severe class imbalance (approximately 1% anomalies), high performance under this metric is critical to minimise manual inspection by domain experts.
-Figure 4 shows representative receiver operation characteristic and Precision-Recall curves for the hourglass anomaly class after the final active learning cycle.
-Our approach achieved an AUROC of $0.96$ and AUPRC of $0.80$, indicating strong separation between nominal and anomalous samples.
-Figure 4: Receiver operation characteristic (left) and Precision-Recall (right) curves after three active learning cycles for the miniImageNet ‘hourglass’ anomaly class starting from five labelled anomalies, adding ten after each iteration.
+We deliberately selected five distinct object classes as anomalies to evaluate our model’s performance across diverse visual features: Guitar, Hourglass, Printer, Piano, and Orange.
+Initial labelled data included 500 images (5 anomalies, 495 normal).
+Each training step utilised an additional randomly sampled pool of 10,000 unlabelled images.
+We summarise our experimental results on the miniImageNet dataset, focusing initially on the class Hourglass as a representative example.
+Results for the other evaluated anomaly classes (Guitar, Printer, Piano, and Orange) are presented in Table 1.
+Given the severe class imbalance (approximately 1% anomalies), high performance under the efficiency metric is critical to minimise manual inspection by domain experts.
+Figure 4 shows the Receiver Operating Characteristic (ROC) curve – plotting the TPR against the FPR – and the Precision-Recall (PR) curve for the Hourglass anomaly class after the final active learning cycle.
+Our approach achieved an AUROC of $0.96$ and AUPRC of $0.83$, indicating strong separation between normal and anomalous samples.
+Figure 4: ROC (left) and PR (right) curves after three active learning cycles for the miniImageNet Hourglass anomaly class starting from five labelled anomalies, adding ten after each cycle.
 The model achieves an AUROC of 0.96 and an AUPRC of 0.80, highlighting robust anomaly detection capability under severe class imbalance.
-Figure 5 shows the anomaly detection efficiency curve after three iterations for the hourglass anomaly class.
-Remarkably, the method identifies 74.7% of anomalies within the top 1% highest-scoring images, demonstrating very high capability to retrieve relevant anomalies for the user.
+Figure 5 shows the Anomaly Detection Efficiency curve after three cycles for the Hourglass anomaly class.
+The method identifies 74.7% of anomalies within the top 1% highest-scoring images, demonstrating high capability to retrieve the specified anomaly for the user.
 The top-scoring 0.1% data samples are all correctly identified anomalies.
 Successive active learning only marginally increase performance.
-A likely reason seems that learning gains from the initial sampling data, which remain the majority of labelled data throughout, can already be achieved within the first iteration.
-(a) Efficiency progression across iterations: efficiency improves quickly, with strong performance already after the first iteration. The red dashed line represents perfect detection performance.
-(b) Final performance after three iterations: 74.7% of all anomalies are found within the top 1% of the ranked predictions, with 10.3% already recovered in the top 0.1%.
+A likely reason is that learning gains from the initial sampling data, which remain the majority of labelled data throughout, are already achieved within the first cycle.
+(a) Anomaly Detection Efficiency progression across cycles: efficiency improves quickly, with strong performance already after the first cycle.
+The red dashed line represents perfect detection performance.
+(b) Final performance after three cycles: 74.7% of all anomalies are found within the top 1% of the ranked predictions, with 10.3% already recovered in the top 0.1%.
 Figure 5: Anomaly Detection Efficiency curves for the Hourglass class in miniImageNet.
@@
 The y-axis shows the percentage of true anomalies recovered within the inspected data subset.
-Active learning started with five labelled anomalies and adds ten samples after each iteration.
-Table 4.2 summarizes performance metrics across all tested anomaly classes.
-The model consistently achieves high AUROC and AUPRC values, identifying approximately three-quarters of anomalies in the top 1% scoring images, and attaining precision between 65% and 81%.
-Average precisions for the top-scoring 0.1% of the data were 99.7%.
+Active learning started with five labelled anomalies and adds ten samples after each cycle.
+Table 1 summarises performance metrics across all tested anomaly classes.
+The model consistently achieves high AUROC and AUPRC values, identifying approximately three-quarters of anomalies in the top 1% scoring images, and attaining precision between 73% and 81%.
+Average precisions for the top-scoring 0.1% of the data were 100%.
 This indicates strong practical utility for human reviewers who would want to quickly identify additional anomalies.
-Table 1: Evaluation on unlabelled data for miniImageNet with different classes used as anomaly after three iterations with active learning.
+Table 1: Evaluation on unlabelled data for miniImageNet with different classes used as anomaly after three cycles with active learning.
 At 100% precision no more anomalies could have been found in the top-scoring 0.1% or 1% of the dataset, respectively.
-Training started from five labelled anomalies, adding ten after each iteration.
-[…]
-=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Anomaly Class AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
-Guitar 0.95 0.76 10.24 100.00 72.32 70.19
-Hourglass 0.96 0.80 10.24 100.00 74.72 72.52
-Printer 0.98 0.86 10.24 100.00 81.44 79.04
-Piano 0.95 0.75 10.24 100.00 71.36 69.25
-Orange 0.92 0.65 10.24 100.00 64.16 62.27
-Mean 0.95 $\pm$ 0.02 0.76 $\pm$ 0.08 10.24 $\pm$ 0.00 100.00 $\pm$ 0.00 72.80 $\pm$ 6.23 70.65 $\pm$ 6.05
+Training started from five labelled anomalies, adding ten after each cycle.
+Anomaly Class AUROC AUPRC
+Precision in
+top-scoring 0.1% [%]
+Precision in
+top-scoring 1% [%]
+Guitar 0.95 0.75 100.00 68.63
+Hourglass 0.97 0.84 100.00 76.40
+Printer 0.97 0.85 100.00 79.35
+Piano 0.97 0.83 100.00 78.26
+Orange 0.95 0.82 100.00 76.55
+Mean 0.96 $\pm$ 0.01 0.82 $\pm$ 0.04 100.00 $\pm$ 0.00 75.84 $\pm$ 4.21
 In summary, these results demonstrate the effectiveness of our approach, highlighting its practical value by consistently ranking most anomalies highly and significantly reducing the labelling workload required from human experts for enabling such a result by starting from only five labelled samples.
-4.3 GalaxyMNIST Results
-For GalaxyMNIST, we illustrate our results using the Unbarred Spiral class as representative anomalies from the dataset.
-Notably, in this case the fraction of anomalies in the dataset (25%) is much higher compared to miniImageNet (1%).
-Performance Metrics and Efficiency
-Receiver operation characteristic and Precision-Recall curves (Figure 6) confirm the robustness of the model.
-Despite GalaxyMNIST’s classes being visually similar and challenging to distinguish even for human observers, the model achieves strong discrimination performance (AUROC: 0.889, AUPRC: 0.774), significantly above random chance.
-Figure 6: Receiver operation characteristic (left) and Precision-Recall (right) curves for the Unbarred Spiral anomaly class from GalaxyMNIST.
+GalaxyMNIST (Walmsley et al., 2022) comprises annotated galaxy images derived from the Galaxy Zoo citizen science project (Lintott et al., 2008), specifically the Galaxy Zoo: DECaLS campaign (Walmsley et al., 2022).
+Images originally have a resolution of $64\times 64$ pixels, which are upsampled to $224\times 224$ pixels for consistency across experiments.
+The dataset contains four morphological classes annotated by volunteers: Smooth Round, Smooth Cigar-shaped, Edge-on Disk, and Unbarred Spiral.
+For evaluation, we treated each morphological type sequentially as anomalies, using the other three classes as normal samples.
+We started experiments with a small labelled set of 40 images (10 anomalies, 30 normal).
+Notably, the fraction of anomalies in this dataset (25%) is much higher compared to miniImageNet (1%).
+As before, each training step utilised an additional randomly sampled pool of 10,000 unlabelled images.
+Figure 6 illustrates our results for the GalaxyMNIST dataset using the Unbarred Spiral class as representative anomalies.
+ROC and PR curves for this class confirm the robustness of the model.
+Despite GalaxyMNIST’s classes being visually similar and challenging to distinguish even for human observers, the model achieves excellent performance metrics (AUROC: 0.89, AUPRC: 0.77), significantly above random chance.
+Figure 6: ROC (left) and PR (right) curves for the Unbarred Spiral anomaly class from GalaxyMNIST.
 High AUROC and AUPRC indicate effective anomaly identification despite the inherent morphological similarities across galaxy classes.
-Training started from ten labelled anomalies, adding ten after each iteration.
-Figure 7 shows how anomaly detection efficiency evolves across iterations.
-A clear improvement of the model is visible after each active learning step, indicating significantly enhanced anomaly ranking capabilities.
-(a) Active learning efficiency across iterations: progressive improvements are visible over three iterations, with more anomalies recovered early in the ranked predictions.
-(b) Final performance after three iterations: within the top 1% of the ranked predictions only anomalies are found
+Training started from ten labelled anomalies, adding ten after each cycle.
+Figure 7 a) shows how Anomaly Detection Efficiency evolves across cycles.
+A clear improvement of the model is visible after each active learning step.
+The final Anomaly Detection Efficiency curve in Figure 7 b) demonstrates strong final performance, particularly at lower inspection percentages.
+Remarkably, within the top 0.1% highest-scoring samples, 100% precision was achieved, thus successfully identifying anomalies immediately within a very small subset of the data.
+Successive active learning continue to increase performance in contrast to the experiments on miniImageNet.
+One potential reason may lie in the overall greater number of anomalous samples in the dataset, leading to a potentially higher heterogeneity of anomalies.
+(a) Active learning efficiency across cycles: progressive improvements are visible over three cycles, with more anomalies recovered early in the ranked predictions.
+(b) Final performance after three cycles: within the top 1% of the ranked predictions, only anomalies are found.
 Figure 7: Anomaly Detection Efficiency curves for the Unbarred Spiral class in GalaxyMNIST.
@@
 The model is trained using ten labelled anomalies initially, with ten more added after each active learning cycle.
-The final anomaly detection efficiency curve in Figure 7 demonstrates strong final performance, particularly at lower inspection percentages.
-Remarkably, within the top 0.1% highest-scoring samples, 100% precision was achieved, thus successfully identifying anomalies immediately within a very small subset of the data.
-Successive active learning continue to increase performance in contrast to the experiments on MiniImageNet.
-One potential reason may lie in the overall greater number of anomalous samples in the dataset, leading to a potentially higher heterogeneity of anomalies.
-Table 4.3 summarises results across all anomaly classes of GalaxyMNIST.
-Table 2: Evaluation on unlabelled data for GalaxyMNIST with different classes used as anomaly after three iterations with active learning.
-At 100% precision no more anomalies could have been found in the top-scoring 0.1% or 1% of the dataset, respectively.
-Training started from ten labelled anomalies, adding ten after each iteration.
-[…]
-=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Anomaly Class AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
-Smooth Round 0.97 0.91 0.36 100.00 3.97 98.99
-Smooth Cigar-shaped 0.75 0.51 0.28 77.78 3.36 83.84
-Edge-on Disk 0.85 0.66 0.36 100.00 3.72 92.93
-Unbarred Spiral 0.89 0.77 0.36 100.00 3.93 97.98
-Mean 0.86 $\pm$ 0.09 0.71 $\pm$ 0.17 0.34 $\pm$ 0.04 94.44 $\pm$ 11.11 3.75 $\pm$ 0.28 93.44 $\pm$ 6.92
+Table 2 summarises results across all GalaxyMNIST anomaly classes.
 Overall, anomaly detection performance on GalaxyMNIST varied between classes, reflecting visual and morphological similarities inherent in the dataset.
@@
 Conversely, classes like Smooth Cigar-shaped and Edge-on Disk led to lower AUROC and AUPRC values, highlighting the inherent difficulty and potential ambiguity in visually separating these morphologies.
-4.4 Ablation Studies
-To evaluate the robustness and practical sensitivity of our approach to key training parameters, we conducted targeted ablation studies on the miniImageNet dataset, specifically focusing on the class guitar as anomalies.
-This class was picked as we achieved median performance on it.
-4.4.1 Number of Initial Labels
+Table 2: Evaluation on unlabelled data for GalaxyMNIST with different classes used as anomaly after three cycles with active learning.
+At 100% precision no more anomalies could have been found in the top-scoring 0.1% or 1% of the dataset, respectively.
+Training started from ten labelled anomalies, adding ten after each cycle.
+Anomaly Class AUROC AUPRC
+Precision in
+top-scoring 0.1% [%]
+Precision in
+top-scoring 1% [%]
+Smooth Round 0.98 0.95 100.00 100.00
+Smooth Cigar-shaped 0.74 0.51 88.89 78.79
+Edge-on Disk 0.89 0.74 88.89 96.97
+Unbarred Spiral 0.94 0.89 100.00 100.00
+Mean 0.89 $\pm$ 0.10 0.77 $\pm$ 0.20 94.44 $\pm$ 6.41 93.94 $\pm$ 10.20
+3.3 Galaxy Zoo - The Galaxy Challenge
+We also benchmark AnomalyMatch on the ’Galaxy Zoo - The Galaxy Challenge’ dataset, which is a subset of $\sim$60,000 galaxy images extracted from Galaxy Zoo 2 (Willett et al., 2013) for the ’Galaxy Challenge’ Kaggle competition.
+The dataset consists of images of galaxies where citizen scientists classified the galaxy morphology.
+One of the classifications they made of importance to us was whether the central source was unusual or ’odd’.
+Each galaxy was seen by at least 40 citizen scientists.
+Each object is ranked by the percentage of users which classified this image accordingly.
+This test set is chosen to directly compare to the results from Lochner & Bassett (2021), who tested the effectiveness of Astronomaly at identifying ’odd’ galaxies in this dataset.
+We follow their approach and define an anomalous galaxy as one for which 90% of the volunteers classified it as ’odd’.
+For training, we randomly sampled (with a fixed seed) three distinct subsets from the overall Galaxy Challenge dataset:
+  • 100 normal and 100 anomalous objects;
+  • 200 normal and 200 anomalous objects;
+  • 394 normal and 6 anomalous objects.
+After each training cycle, we correct the top 10 highest-scoring mislabelled normal and confirm the top 10 highest-scoring anomalous objects.
+To achieve robust measurements of our performance, we run these different configurations with 1 well used seed and 8 random seeds [^5] The following seeds were used, and no further seeds were run: 42, 76032, 730, 83209, 13798, 4538, 5923, 99271, 3762 and average their results.
+This provides us with better estimates of performance and the variation in our method over different selections of training data, and different selections of active learning sets that are informing the model.
+Figure 8 shows the number of anomalies identified in the top N-ranked images based on the score given by AnomalyMatch using a training set of 200 nominal and 200 anomaly images (bringing the total to 440 after active learning).
+The rank here is a simple ordering of the dataset from most anomalous to least anomalous by score, where objects with identical score will be randomised in their ordering.
+As stated previously, we have a soft truth label of what is and is not an anomaly by taking the 90% voter fraction cut used in Lochner & Bassett (2021).
+We find that, overall, our performance is similar to that of Astronomaly.
+At high ranks (low index), our performance is below that of Astronomaly while at lower rank (increasing index) AnomalyMatch begins to outperform it.
+However, both methods are below a perfect 1:1 detection.
+To investigate the cause of this low performance, we vary the definition of an anomaly by applying different voter fraction (GZ Score) cutoffs.
+We find that the performance is sensitive to the chosen cutoff.
+At 0.90 we are searching for 704 anomalies (not including the 220 used in training).
+Altering this voter cutoff affects our performance as the number of anomalies we are attempting to identify changes.
+For instance, with a voter cutoff of 0.8 ($N_{\text{anomalies}}=2,058$) we detect 372 anomalies ($\sim 20\%$) in the top 1,000 scored images.
+Similarly, for 0.7 ($N_{\text{anomalies}}=3,670$) we detect 490 anomalies ($\sim 15\%$ in the top 1,000 scores.
+Making the cutoff more stringent, to 0.95, reduces the number to search for to $N_{\text{anomalies}}=154$, where we detect 109 ($\sim 70\%$) in the top 1,000 ranked images.
+This shows the volatility in the definition of an anomaly in this dataset.
+Furthermore, we cannot directly compare this to Astronomaly, as they do not investigate the variance in their model performance nor the effect of changing the voter cutoff.
+To explore the agreement or disagreement between the volunteers and AnomalyMatch, we investigate differences in ranking given by each.
+We subtract the citizen scientist rank from the rank found by AnomalyMatch.
+If the Galaxy Zoo rank and AnomalyMatch rank are in agreement, this difference is close to zero.
+If the citizen scientists and AnomalyMatch disagree, this rank difference is large.
+Figure 9 shows the distribution for the entire sample (blue) and only the objects defined as anomalies (red).
+We find that the volunteers and AnomalyMatch are in close agreement (rank difference $\leq 1000$) for 273 ($\sim$40%) of anomalies.
+Figure 10 shows a grid of example sources ordered by agreement between AnomalyMatch scores and GZ Scores.
+Objects are sorted from left to right according to the anomaly score assigned by AnomalyMatch.
+Vertically, the objects are split into normal (top) and anomalies (bottom).
+The three rows in each group correspond (from top to bottom) to minimal, median and highest rank disagreement.
+The difference between the volunteer and AnomalyMatch ranks are shown in the inset of each image.
+For the anomalies (red borders) we find that the disagreement can, for example, come from dual core objects, as well as mergers showing limited tidal features.
+Objects scored highly by AnomlyMatch, in general, are often those with clear tidal features and ongoing interactions.
+There is a significant difference in volunteer votes and AnomalyMatch scores for images with colour errors, images containing artifacts and ring galaxies, to which AnomalyMatch assigned low scores.
+These were likely not represented in our training sample and not included while conducting active learning.
+The bottom row of the normal images in Figure 10 shows that AnomalyMatch disagreed with volunteers by ranking as anomalous noisy images or ones with small and close, possibly interacting, companions and weak but possible tidal features.
+Meanwhile, anomaly images of overlapping sources were typically ranked much lower (bottom left) than by users.
+The disagreement in the normal class likely stems from the difficulty for the volunteers to visually identify merger signatures that are more easily found by AnomalyMatch.
+Figure 11 shows the distribution of scores given by AnomalyMatch to normal and anomalous images as ranked by the volunteers.
+The score distribution for normal images (blue) is highly skewed towards low scores.
+In contrast, the distribution for anomalies (red) is comparatively more bimodal.
+While most anomalies recieve high scores, there is also a small overdensity at low scores.
+Examples of anomalies that fall into this overdensity - i.e.
+anomalies ‘missed’ by AnomalyMatch - are the dual-core galaxies in the bottom left of Figure 10.
+This also reinforces that during active learning a user should also consider what is being misclassified by the model to correct its behaviour.
+Table 3 summarises our performance metrics across different breakdowns of labelled vs unlabelled data in training.
+The best performance is achieved by the evenly split model with 200-200 anomaly-normal labels, with top-scoring 0.1% precision of $(40\pm 10)\%$.
+For the more likely real-world setup of 6-394, we measure mean precisions of $(31.93\pm 9.84)$% and $(19.27\pm 3.69)$% in the top-scoring 0.1% and 1% of anomalies observed.
+Table 3: Evaluation on unlabelled data for the Galaxy Zoo 2 dataset with different initial configurations, varying in the number of labels and the proportion of anomalies.
+Averaging results from 9 random seeds.
+Labels
+[Total (Anomalies / Normal)]
+AUROC AUPRC
+Precision in
+top-scoring 0.1% [%]
+Precision in
+top-scoring 1% [%]
+200 (100 / 100) 0.87 $\pm$ 0.02 0.12 $\pm$ 0.01 32.60 20.41
+400 (200 / 200) 0.91 $\pm$ 0.01 0.17 $\pm$ 0.02 39.52 25.00
+400 (6 / 394) 0.71 $\pm$ 0.04 0.06 $\pm$ 0.03 23.67 12.40
+Mean 0.83 $\pm$ 0.02 0.12 $\pm$ 0.01 31.93 $\pm$ 9.84 19.27 $\pm$ 9.18
+Figure 8: Direct comparison of AnomalyMatch to Astronomaly using an initial training set of 200 normal and 200 anomalous label images and adding 20 labels per active learning cycle in three training cycles, for a total of 440 labels.
+For AnomalyMatch, we plot the mean curves over 9 random seeds, with shaded regions indicating the standard deviation.
+The curve for Astronomaly (dashed red) is taken from Figure 5a of Lochner & Bassett (2021).
+A line representing perfect detection (i.e.
+all anomalies are in the top 2,000 ranked indexes) is shown in gray.
+We find that the performance of Astronomaly and AnomalyMatch is comparable, although AnomalyMatch performs slightly worse at initial ranks but then starts to outperform Astronomaly.
+Lowering the voter fraction threshold leads to an improved performance of AnomalyMatch, indicating a high degree of class confusion.
+Figure 9: Comparison of the ranks based on GZ scores and scores from AnomalyMatch for fixed seed 4538 after cycle 3 with the 200 - 200 balanced labelled set.
+It is evident that the ranks assigned by AnomalyMatch are relatively close to the user ranks, while uninteresting objects have a much broader distribution.
+Figure 10: Comparison of the GZ ranks and the AnomalyMatch scores with fixed seed 4538 after cycle 3 for the 200 - 200 balanced labelled set.
+Each column represents a percentile of AnomalyMatch scores (class based) and displays 3 samples (closest -, median - and furthest absolute rank deviation) for normal and anomaly images.
+Positive rank differences indicate that AnomalyMatch gave the object a higher score.
+In the anomalies it can be seen, that objects showing strong tidal features were typically ranked/scored high and closer to the GZ rank, while overlapping galaxies or galaxies with 2 cores were ranked much lower.
+Of the normal images noisy galaxies showing colour gradients were often ranked much higher than in GZ 2.
+In a real world use case a user would recognise such clear mislabelling and amend the dataset accordingly.
+Figure 11: Histogram of the scores computed by AnomalyMatch with a specific seed after cycle 3 with the 200 - 200 balanced labelled set.
+A very clear separation between anomalies and the normal images is present.
+3.4 Number of Initial Labels
 We investigated the impact of varying the number of initial labels while maintaining the proportion of anomalies within these labels.
-Three scenarios were tested: 100 initial labels (one anomaly, 99 nominal), 500 initial labels (five anomalies, 495 nominal), and 1000 initial labels (10 anomalies, 990 nominal).
+Three scenarios were tested: 100 initial labels (one anomaly, 99 normal), 500 initial labels (five anomalies, 495 normal), and 1000 initial labels (10 anomalies, 990 normal).
 After each active learning cycle, we correspondingly increased the number of actively labelled samples per iteration proportionally to the initial sample size: 10 additional labels per iteration for the smallest set, 20 labels for the intermediate set, and 40 labels for the largest set.
-Detailed results for these ablations are summarised in Table 4.4.1.
+Detailed results for these ablations are summarised in Table 4.
+Interestingly, we observed limited differences in performance metrics across all tested initial labelling conditions.
+AUROC values varied by less than $0.03$, and Anomaly Detection Efficiency remained consistently high.
+Notably, all experiments achieved perfect precision at the stringent threshold of inspecting just the top 0.1% scoring data points, further underscoring the robustness and consistency of our active learning pipeline.
+This indicates that our approach effectively utilises available labels and does not significantly benefit from substantially larger initial label sets, highlighting its practical efficacy even with minimal labelled data.
 Table 4: Effect of size of initial labelled dataset on anomaly detection performance in miniImageNet.
-Each configuration starts with a different number of labelled samples and adds different number of labels per iteration.
+Each configuration starts with a different number of labelled samples and adds different number of labels per cycle.
 Despite large differences in initial label counts, performance remains robust across metrics, with perfect precision at the top 0.1% most anomalous samples for all scenarios.
-Beyond 500 (+20) labels no performance gains are observed.
-[…]
-=1.2mm {tabu} | X[0.45cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Labels AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
-100 (1 / 99) + 10 per iteration 0.93 0.65 10.02 100.00 61.66 60.80
-500 (5 / 495) + 20 per iteration 0.96 0.79 10.24 100.00 75.20 72.98
-1000 (10 / 990) + 40 per iteration 0.95 0.78 10.50 100.00 75.00 70.42
-Mean 0.95 $\pm$ 0.02 0.74 $\pm$ 0.08 10.25 $\pm$ 0.24 100.00 $\pm$ 0.00 70.62 $\pm$ 7.76 68.07 $\pm$ 6.42
-Interestingly, we observed limited differences in performance metrics across all tested initial labelling conditions.
-AUROC and AUPRC values varied by less than $0.03$, and anomaly detection efficiency remained consistently high.
-Notably, all experiments achieved perfect precision at the stringent threshold of inspecting just the top 0.1% scoring data points, further underscoring the robustness and consistency of our active learning pipeline.
-This indicates our approach effectively utilises available labels and does not significantly benefit from substantially larger initial label sets, highlighting its practical efficacy even with minimal labelled data.
-4.4.2 Training Iterations
-To assess how training duration per cycle affects anomaly detection, we varied the number of FixMatch training iterations per cycle across four settings: 50, 100, 250, and 500.
-Results are presented in Table 4.4.2.
-Table 4: Impact of FixMatch training duration per active learning step.
-Performance peaks at shorter training cycles (50–100 iterations), with diminishing returns and mild overfitting at longer durations.
-This supports our design choice of brief, iterative updates with rapid model adaptation.
-[…]
-=1.2mm {tabu} | X[0.4cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Iterations AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
-50 0.97 0.77 10.24 100.00 71.52 69.41
-100 0.97 0.79 10.24 100.00 74.72 72.52
-250 0.95 0.78 10.24 100.00 74.08 71.89
-500 0.94 0.77 10.24 100.00 73.44 71.27
-Mean 0.96 $\pm$ 0.02 0.78 $\pm$ 0.01 10.24 $\pm$ 0.00 100.00 $\pm$ 0.00 73.44 $\pm$ 1.38 71.27 $\pm$ 1.34
+Beyond 500 (+20) labels only minimal performance gains are observed.
+Labels
+[Total (Anomalies / Normal)]
+AUROC AUPRC
+Precision in
+top-scoring 0.1% [%]
+Precision in
+top-scoring 1% [%]
+100 (1 / 99) + 10 per iter. 0.92 0.35 59.38 41.82
+500 (5 / 495) + 20 per iter. 0.95 0.76 100.00 69.88
+1000 (10 / 990) + 40 per iter 0.96 0.80 100.00 71.52
+Mean 0.94 $\pm$ 0.02 0.64 $\pm$ 0.25 86.46 $\pm$ 23.45 61.07 $\pm$ 16.69
+3.5 Training Iterations
+To test for the optimal number of iterations in our setup, we varied the number of FixMatch training iterations per cycle across four settings: 50, 100, 250, and 500.
+Table 5 shows the results of these experiments.
 We observe that performance peaks at around 100 training iterations per cycle.
 Using more iterations results in slightly worse AUROC and AUPRC, which we attribute to overfitting on the initially provided labelled data.
-As the active learning loop dynamically updates the training set, excessively long training phases may entrench early biases.
-This supports our design choice of short training cycles, encouraging the model to adapt quickly to newly labelled examples.
-4.4.3 Impact of Active Learning
-We specifically investigated the impact of incorporating active learning by comparing our standard iterative labelling approach against two pure semi-supervised scenarios without additional human-in-the-loop labels, one with the same amount of initial labels and, respectively, same final amount (i.e.
-total number including those added during active learning).
-Figure 8 illustrates anomaly detection efficiency for those scenarios.
-(a) Without active learning (same starting samples)
-(b) Without active learning (same total samples)
-(c) With active learning
-(d) Comparison of approaches
-Figure 8: Anomaly detection efficiency comparison with and without active learning.
-The actively learned model achieves notably better efficiency within the crucial 0.1–1% data inspection range for the same starting labels (compare (a) and (c)), indicating a meaningful advantage in human-in-the-loop anomaly searches.
-However, training with the full set of labels (as would be identified during active learning) from scratch leads to the best performance (b), highlighting the potential of re-using found anomalies in new semi-supervised runs.
-The direct comparison in (d) further illustrates the relative performance across all approaches, demonstrating how active learning bridges the gap between limited-label and full-label scenarios.
-Table 5: Comparison of anomaly detection performance with and without active learning.
-Active learning at same starting labels yields higher AUPRC and precision within the top 1% most anomalous predictions.
-Using all actively gathered labels for semi-supervised training yields the highest performance, suggesting benefits from iterative human-in-the-loop feedback and re-use of confirmed anomalies.
-[…]
-=1.2mm {tabu} | X[0.8cm] | X[0.5cm] | X[0.5cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | X[0.7cm] | Active Learning AUROC AUPRC Anomalies in top-scoring 0.1% [%] Precision in top-scoring 0.1% [%] Anomalies in top-scoring 1% [%] Precision in top-scoring 1% [%]
-Active Learning 0.96 0.79 10.24 100.00 74.56 72.36
-No Active Learning (same starting samples) 0.96 0.70 9.77 98.44 65.89 65.89
-No Active Learning (same total samples) 0.99 0.85 10.41 100.00 78.70 75.16
-Mean 0.97 $\pm$ 0.02 0.78 $\pm$ 0.08 10.14 $\pm$ 0.33 99.48 $\pm$ 0.90 73.05 $\pm$ 6.54 71.14 $\pm$ 4.75
-While passive training (no additional active learning) yielded similar AUROC and AUPRC metrics, active learning exhibited significantly improved practical efficiency in regard to anomaly detection efficiency.
-Crucially, in the most relevant operational range (inspecting the top 0.1%–1% of scored samples), active learning consistently ranked more anomalies higher for the same starting samples.
-For instance, active learning allowed users to detect 75% of anomalies within the top 1% of the ranked dataset, while the passive model lagged behind in top-scoring predictions’ precision (particularly visible in the 0.1% inspection range).
-This highlights a critical trade-off: despite marginally lower overall metric scores, the actively updated model significantly reduces human inspection efforts by focusing the user’s attention effectively, thus strongly supporting the use of active learning in real-world applications.
-However, retraining with all found anomalies still yields the best performance, likely due to a more balanced representation of anomaly samples during training.
-4.5 Discussion
+As the active learning loop dynamically updates the training set, excessively long training phases may lead to overfitting on initial samples.
+Table 5: Impact of FixMatch training duration per active learning step.
+Performance peaks at shorter training cycles (50–100 iterations), with diminishing returns and overfitting at longer durations.
+This supports our design choice of brief, iterative updates with rapid model adaptation.
+Iterations AUROC AUPRC
+Precision in
+top-scoring 0.1% [%]
+Precision in
+top-scoring 1% [%]
+50 0.98 0.81 100.00 75.31
+100 0.95 0.76 100.00 69.72
+250 0.92 0.73 100.00 67.55
+500 0.93 0.73 100.00 67.70
+Mean 0.95 $\pm$ 0.03 0.76 $\pm$ 0.04 100.00 $\pm$ 0.00 70.07 $\pm$ 3.63
+4 Discussion
 Our results highlight the practical efficiency and robustness of the AnomalyMatch framework, demonstrating significant potential to substantially reduce the manual review burden for anomaly identification tasks.
-The method efficiently leverages both computational resources and human labelling effort, achieving strong performance even with very limited labelled data.
+The method efficiently leverages both computational resources and human labelling effort, achieving strong performance even with limited labelled data.
 Crucially, our experiments reveal diminishing returns from significantly increasing the number of initial labels.
-This effect is consistent with earlier findings in semi-supervised learning theory and empirical studies that suggest performance gains saturate quickly beyond a modest label count [30].
-This outcome underscores the method’s effective utilisation of limited supervision, suggesting that label sparsity — a common real-world constraint — is not a major limitation in practice.
-The integration of active learning notably enhances anomaly detection efficiency, particularly in the critical inspection range (top 0.1%–1% of ranked data).
-While active learning’s impact on global performance metrics such as AUROC and AUPRC might appear moderate, its practical value in real-world scenarios cannot be overstated, as domain experts typically inspect only a small fraction of data.
-Thus, actively guided iterative labelling can substantially increase the productivity of domain experts, aligning well with real-world needs.
+This effect is consistent with earlier findings in semi-supervised methodologies and empirical studies that suggest performance gains saturate quickly beyond a modest label count (Oliver et al., 2018).
+This outcome underscores the method’s effective utilisation of limited supervision, suggesting that label sparsity – a common real-world constraint – is not a major limitation in practice.
 A key finding of our study is the pronounced risk of overfitting when extending training beyond the optimal number of iterations per active learning cycle.
-Notably, our experiments suggest the ideal balance is around 100 iterations per active learning cycle; longer cycles appear to reduce overall anomaly detection efficiency, highlighting the need for careful tuning and early stopping strategies.
-Such overfitting risk in SSL has been previously observed, especially in low-label regimes without strong regularisation [31].
+Notably, these suggest the ideal balance is around 100 iterations per active learning cycle; longer cycles appear to reduce overall Anomaly Detection Efficiency, highlighting the need for careful tuning and early stopping strategies.
+Such overfitting risk in semi-supervised learning has been previously observed, especially in low-label regimes without strong regularisation (Arazo et al., 2020).
 Given that original FixMatch and its derivatives such as MSMatch typically employ extensive training iterations, our adaptation of shorter, frequent retraining cycles emerges as a beneficial modification tailored to active learning scenarios with heavily imbalanced classes and a simplified binary classification task.
-Additionally, while we have demonstrated effectiveness across datasets with drastically different class imbalance characteristics –— extremely skewed (miniImageNet) and moderately balanced (GalaxyMNIST) —– further research is necessary to investigate the lower bounds of label numbers needed for robust anomaly detection performance.
-Preliminary findings from an accompanying paper focused on the Hubble Space Telescope dataset [10] suggest that AnomalyMatch might remain effective with even fewer initial labels, further reducing labelling costs.
+Further, preliminary findings from an accompanying paper focused on the Hubble Space Telescope dataset (O’Ryan & Gómez, 2025) suggest that AnomalyMatch might remain effective with even fewer initial labels, further reducing labelling costs.
 There, we also utilised a hybrid approach: initially using active learning to identify a sufficient number of anomalies, then transitioning to a fully semi-supervised regime for optimal performance.
+While we have demonstrated effectiveness across datasets with drastically different class imbalance characteristics – extremely skewed (miniImageNet) and moderately balanced (GalaxyMNIST) – further research is necessary to investigate the lower bounds of label numbers needed for robust anomaly detection performance.
+Our performance evaluation on the Galaxy Zoo 2 dataset is more complex.
+To explore potential reasons behind the performance metrics we find, we identify the anomalies that AnomalyMatch scored lowly and the normal images that it scored highly.
+Figure 12 shows such images: on the left are images that the volunteers scored >0.90>0.90 while AnomalyMatch gave a score of <0.2<0.2 (N=112=112), the right shows images AnomalyMatch gave scores >0.95>0.95 and the volunteers scored <0.90<0.90 (N=82=82).
+Figure 12: A sample of 50 images given either high score by AnomalyMatch but a low one in Galaxy Zoo 2 or vice versa.
+Left: AnomalyMatch gave a low score (<0.2<0.2) but the object was classified as an anomaly (Galaxy Zoo score >0.90>0.90).
+Right: AnomalyMatch gave a high score (>0.95>0.95) but this was not classified as an anomaly (Galaxy Zoo score <0.90<0.90).
+The total number of images satisfying both criteria is 194.
+The left of Figure 12 shows anomalies that AnomalyMatch "missed".
+A key example in this representative set is that of ring galaxies.
+It is likely that this class was not represented in active learning.
+In practice, a user could correct these disagreements in the active learning stage and therefore circumvent such misclassifications, improving measured performance.
+The right hand side of Figure 12 shows objects that AnomalyMatch scored >0.95>0.95, but were not classified as anomalies in the Galaxy Zoo 2 dataset.
+We find many ‘odd’ galaxies, such as a polar ring galaxy (last in third row) as well as many overlapping galaxies that could be of interest.
+There are also examples of galaxies overlapped with point sources or containing speckles which.
+It is important to note the Galaxy Zoo 2 scores given to these objects.
+Many objects have a volunteer vote fraction close to 0.90, our chosen threshold for separating nominal from anomalous objects.
+These edge cases may therefore be real anomalies that were classified as nominal, reducing our performance.
+While these examples are drawn from our testing set, it is likely that similar misclassifications exist in the Galaxy Zoo 2 dataset used for training.
+Individual misclassifications in the training set can affect the model’s ability to correctly identify edge cases.
 4.1 Limitations
-Despite strong results, our study acknowledges several important limitations that warrant further exploration:
-Idealised Active Learning Assumptions:
+Despite strong results, our study acknowledges several important limitations that warrant further exploration.
+Addressing these limitations in future work will further strengthen the practical applicability of our approach, enabling its robust deployment in diverse, realistic scientific and industrial contexts.
+Idealised Active Learning Assumptions.
 Our experiments assumed an idealised active learning environment where human annotations are consistently accurate.
 In practical deployments, labelling errors and inconsistencies are likely, particularly in ambiguous cases.
-Additionally, our experiments assumed balanced incremental labelling (equal numbers of anomalies and nominal cases per active learning cycle), which may differ significantly in real-world usage where nominal labels are generally easier and cheaper to obtain than anomalies.
-Pretraining Bias:
+Additionally, our experiments assumed balanced incremental labelling (equal numbers of anomalies and normal cases per active learning cycle), which may differ significantly in real-world usage where normal labels are generally easier and cheaper to obtain than anomalies.
+Pretraining Bias.
 Performance on the miniImageNet dataset likely benefits from the EfficientNet backbone’s ImageNet pretraining.
@@
 Future studies should assess performance sensitivity when employing randomly initialised neural networks or representations pretrained on domain-specific data.
-Limited Exploration of Class Imbalance Scenarios:
+Limited Exploration of Class Imbalance Scenarios.
 While we investigated datasets at two extremes of class imbalance (1% vs.
@@
 Additional research is needed to characterise model behaviour across a broader spectrum of imbalance ratios.
-Anomaly Homogeneity Assumption:
-We simplified the anomaly detection task by treating all anomalies as belonging to a single implicit category distinct from nominal instances.
-This assumption might limit applicability in scenarios with highly heterogeneous anomalies, each requiring distinct learned representations.
+Anomaly Homogeneity Assumption.
+We simplified the anomaly detection task by treating all anomalies as belonging to a single implicit category distinct from normal instances.
+This assumption would limit applicability in scenarios with highly heterogeneous anomalies, each requiring distinct learned representations.
 Incorporating methods explicitly capable of handling diverse anomaly categories could enhance generalisation.
-Impact of Labelling Errors:
+Impact of Labelling Errors.
 Our study does not examine the robustness of AnomalyMatch to labelling inaccuracies, an important consideration given the potential for human annotation errors during active learning.
 Exploring training strategies that explicitly handle or mitigate labelling noise is necessary for real-world robustness.
-Addressing these limitations in future work will further strengthen the practical applicability of our approach, enabling its robust deployment in diverse, realistic scientific and industrial contexts.
-5 Conclusion and Outlook
+Variation in Model Seeds In this study, we have only conducted a very minimal investigation into the different seeds and hyper-parameters of the model we use.
+Figure 8 shows the standard deviation in our performance when comparing an ensemble of AnomalyMatch models to Astronomaly.
+As shown there, performance may vary on the anomalies detected dependent on the random seed used.
+This is not unusual in the context of using neural networks.
+Further studies are required to fully investigate the effects of hyper parameters and seeds on the performance of the model, however, for now we can state the model performance is comparable to tested classical methods such as Astronomaly.
+5 Conclusion
 In this study, we introduced AnomalyMatch, a robust semi-supervised active learning framework tailored for efficiently identifying anomalies in large-scale image datasets.
-Our pipeline, combining FixMatch-based semi-supervised learning, EfficientNet as a computational backbone, and an interactive graphical user interface, demonstrated excellent performance across diverse datasets (GalaxyMNIST and miniImageNet), highlighting its capability to effectively operate under severe class imbalance and label scarcity—conditions frequently encountered in astronomy.
+Our pipeline, combining FixMatch-based semi-supervised learning, EfficientNet as a computational backbone, and an interactive GUI, demonstrated excellent performance across diverse datasets (GalaxyMNIST and miniImageNet), highlighting its capability to effectively operate under severe class imbalance and label scarcity – conditions frequently encountered in astronomy.
 By significantly reducing the manual review burden, AnomalyMatch offers substantial potential to streamline anomaly discovery processes, effectively harnessing minimal human input and maximising the efficiency of scarce domain-expert resources.
+The results presented in this work, along with those in the companion paper (O’Ryan & Gómez, 2025), underscore the versatility, broad applicability, and potential of AnomalyMatch.
+Its ability to handle diverse data formats commonly used in astronomy, combined with a user-friendly GUI and integration within ESA Datalabs – granting direct access to repositories such as the Euclid data volumes – make it a powerful tool for identifying rare astronomical objects in the era of increasingly data-rich sky surveys.
 Looking ahead, several promising opportunities exist to further refine and extend the capabilities of AnomalyMatch.
-Incorporating explainable AI methods, such as attention maps, SHAP [32] values, or feature attribution techniques, could provide crucial interpretability and insight into model decision-making, increasing trust and facilitating model validation by domain experts.
-Additionally, extending the model to handle multimodal data—such as integrating image information with complementary spectral or time-series data—would allow richer contextual analysis, potentially enhancing anomaly detection accuracy in complex scientific scenarios.
+Incorporating explainable AI methods – such as attention maps, SHapley Additive exPlanations (SHAP; Lundberg & Lee, 2017) values, or feature attribution techniques – could provide crucial interpretability and insight into model decision-making, increasing trust and facilitating model validation by domain experts.
+Additionally, extending the model to handle multimodal data – e.g., integrating image information with complementary spectral or time-series data – would allow richer contextual analysis, potentially enhancing anomaly detection accuracy in complex scientific scenarios.
 Further improvements could come from developing adaptive training strategies that dynamically calibrate training steps and parameters in response to real-time performance metrics, thus mitigating overfitting risks identified in our current approach.
@@
 Pursuing these enhancements will solidify AnomalyMatch as a versatile and impactful framework, beneficial not only to astronomy but broadly to scientific domains confronted with similar data challenges.
+Data availability statement
+All data used in benchmarking is readily available from the websites listed below.
+The AnomalyMatch algorithm is available on GitHub at https://github.com/esa/AnomalyMatch and on ESA Datalabs at https://datalabs.esa.int/.
+The sources of benchmarking data are:
+  1. 1.
+    The Galaxy Zoo sample is available on Kaggle in the Galaxy Zoo - The Galaxy Challenge.
+All images and labels are available at https://www.kaggle.com/c/galaxy-zoo-the-galaxy-challenge.
+This dataset is described in Willett et al.
+(2013) and Lochner & Bassett (2021).
+  2. 2.
+    The GalaxyMNIST dataset is available from https://github.com/mwalmsley/galaxy_mnist.
+  3. 3.
+    The miniImageNet dataset is available from https://www.kaggle.com/datasets/arjunashok33/miniimagenet.
 Acknowledgments
 This work made use of AI tools for code generation and writing support.
-The authors would like to acknowledge valuable feedback received from Pedro Mas Buitrago, Laslo Erik Ruhberg and Maria Teresa Nardone.
+The authors acknowledge valuable feedback received from Pedro Mas Buitrago.
+DOR acknowledges the support of the European Space Agency Research Fellowship Program.
+This research made use of many open-source Python packages and scientific computing packages.
+These include Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Matplotlib (Hunter, 2007), Pandas (Wes McKinney, 2010), PyTorch (Paszke et al., 2019), Numpy (Harris et al., 2020), scikit-image (Van der Walt et al., 2014) and Pillow (Clark, 2015).
+Appendix A Mathematical formulation of performance metrics
+AUROC and AUPRC are computed based on the counts of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN).
+Let $t\in[0,1]$ be the classification threshold.
+Then, AUROC and AUPRC can be defined as:
+$\text{AUROC}=\int_{0}^{1}\text{TPR}(t)\,d\text{FPR}(t)\ ,$ (2)
+$\text{AUPRC}=\int_{0}^{1}\text{Precision}(t)\,d\text{Recall}(t)\ ;$ (3)
+Where:
+$\text{True Positive Rate}(t)=\text{TPR}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FN}(t)}\ ,$ (4)
+$\text{False Positive Rate}(t)=\text{FPR}(t)=\frac{\text{FP}(t)}{\text{FP}(t)+\text{TN}(t)}\ ,$ (5)
+$\text{Precision}(t)=\frac{\text{TP}(t)}{\text{TP}(t)+\text{FP}(t)}\ ,$ (6)
+$\text{Recall}(t)=\text{TPR}(t)\ .$ (7)
+The Anomaly Detection Efficiency is defined as the fraction of correctly identified anomalies within a specified percentage of top-scoring samples:
+[…]